Jefferson Lab HALL A Logbook for August 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 73269 08/31/01 16:58 jpchen cooling flow; oven/heater; laser hut AC/air flow
  • 73268 08/31/01 16:53 jpchen Shapiro removed and in EEl target lab now.
  • 73267 08/31/01 16:22 sulkosky Parity DAQ-Limit on # of events changed.
  • 73264 08/31/01 16:16 apar Start_of_Run_Parity_3366
  • 73260 08/31/01 15:03 apar Start_of_Run_Parity_3365
  • 73255 08/31/01 13:28 apar Start_of_Run_Parity_3364
  • 73253 08/31/01 09:24 W. Korsch Pedestal Run
  • 73251 08/31/01 08:21 apar Start_of_Run_Parity_3363
  • 73250 08/31/01 08:09 Nilanga Shift summary
  • 73246 08/31/01 07:56 apar Start_of_Run_Parity_3362
  • 73241 08/31/01 07:00 apar Start_of_Run_Parity_3361
  • 73237 08/31/01 06:07 X. Zheng Target - NMR 38.62%, EPR 39.92%
  • 73235 08/31/01 06:04 apar Start_of_Run_Parity_3360
  • 73232 08/31/01 06:00 adaq Runsummary_Left_HRS
  • 73231 08/31/01 06:00 adaq Runsummary_Right_HRS
  • 73227 08/31/01 04:56 apar Start_of_Run_Parity_3359
  • 73223 08/31/01 03:58 apar Start_of_Run_Parity_3358
  • 73217 08/31/01 03:02 apar Start_of_Run_Parity_3357
  • 73212 08/31/01 02:06 apar Start_of_Run_Parity_3356
  • 73207 08/31/01 01:11 apar Start_of_Run_Parity_3355
  • 73202 08/31/01 00:15 apar Start_of_Run_Parity_3354
  • Thursday

  • 73196 08/30/01 23:19 apar Start_of_Run_Parity_3353
  • 73192 08/30/01 22:27 Vince, Xiaof Target NMR: 38.6% EPR: 38.9%
  • 73190 08/30/01 22:24 apar Start_of_Run_Parity_3352
  • 73185 08/30/01 21:02 apar Start_of_Run_Parity_3351
  • 73179 08/30/01 20:07 apar Start_of_Run_Parity_3350
  • 73175 08/30/01 19:12 apar Start_of_Run_Parity_3349
  • 73170 08/30/01 18:17 apar Start_of_Run_Parity_3348
  • 73164 08/30/01 17:21 apar Start_of_Run_Parity_3347
  • 73160 08/30/01 16:32 Sulkosky Target-NMR: 38.7 %
  • 73159 08/30/01 16:15 Xiaofeng, Wo Target EPR: 39.8%
  • 73157 08/30/01 16:13 apar Start_of_Run_Parity_3346
  • 73155 08/30/01 16:04 x. jiang shift summary, day shift
  • 73150 08/30/01 14:46 apar Start_of_Run_Parity_3345
  • 73147 08/30/01 13:58 x.jiang left arm HV screen at run 3246/31246
  • 73146 08/30/01 13:57 x. jiang right-arm HV screem at run 3246/31246
  • 73145 08/30/01 13:54 x.jiang main control screen at run 3246/31246
  • 73143 08/30/01 13:49 apar Start_of_Run_Parity_3344
  • 73138 08/30/01 12:54 apar Start_of_Run_Parity_3343
  • 73133 08/30/01 11:59 apar Start_of_Run_Parity_3342
  • 73128 08/30/01 11:04 apar Start_of_Run_Parity_3341
  • 73123 08/30/01 10:08 x.jiang target rotatetd to 270 degree
  • 73122 08/30/01 10:08 apar Start_of_Run_Parity_3340
  • 73120 08/30/01 10:03 W. Korsch Target -EPR: 40.4%
  • 73116 08/30/01 08:35 apar Start_of_Run_Parity_3339
  • 73111 08/30/01 08:02 A. Deur shift summary
  • 73109 08/30/01 07:39 apar Start_of_Run_Parity_3338
  • 73103 08/30/01 06:44 apar Start_of_Run_Parity_3337
  • 73100 08/30/01 06:00 adaq Runsummary_Left_HRS
  • 73099 08/30/01 06:00 adaq Runsummary_Right_HRS
  • 73097 08/30/01 05:48 apar Start_of_Run_Parity_3336
  • 73093 08/30/01 05:04 X. Zheng Target - NMR 39.50%, EPR 40.19%
  • 73091 08/30/01 04:52 apar Start_of_Run_Parity_3335
  • 73086 08/30/01 03:29 apar Start_of_Run_Parity_3334
  • 73081 08/30/01 02:33 apar Start_of_Run_Parity_3333
  • 73076 08/30/01 01:39 apar Start_of_Run_Parity_3332
  • 73071 08/30/01 00:42 apar Start_of_Run_Parity_3331
  • 73067 08/30/01 00:06 P.Zolnierczu swing shift summary
  • Wednesday

  • 73063 08/29/01 23:46 jpchen/Vince Target: EPR/NMR=41.39%/40.9%
  • 73062 08/29/01 23:46 JPChen/Sulko Target-ref. cell filled with N2.
  • 73060 08/29/01 23:41 Sulkosky TARGET-Field configuration change to 90 degrees.
  • 73057 08/29/01 23:29 jpchen Resume data taking at kin4 out/90
  • 73056 08/29/01 23:28 apar Start_of_Run_Parity_3329
  • 73049 08/29/01 22:39 apar Start_of_Run_Parity_3328
  • 73046 08/29/01 22:33 Eugene Moller results
  • 73042 08/29/01 22:16 apar Start_of_Run_Parity_3327
  • 73041 08/29/01 22:16 jpchen OTR data of 240 um was not valid
  • 73040 08/29/01 22:12 jpchen moved to ref cell
  • 73036 08/29/01 22:04 apar Start_of_Run_Parity_3326
  • 73033 08/29/01 21:56 jpchen OTR spot check
  • 73030 08/29/01 21:54 apar Start_of_Run_Parity_3325
  • 73029 08/29/01 21:42 jpchen Moller completed. BeO in waiting for normal beam back
  • 73028 08/29/01 20:06 Sulkosky Target-Longitudinal pumping at 0 degrees.
  • 73027 08/29/01 19:48 jpchen/Vince target: EPR=39.65%, NMR=39.2%
  • 73024 08/29/01 18:09 apar Start_of_Run_Parity_3324
  • 73019 08/29/01 17:10 apar Start_of_Run_Parity_3323
  • 73015 08/29/01 16:38 Zolnierczuk EPR=39.71%
  • 73013 08/29/01 16:14 apar Start_of_Run_Parity_3322
  • 73011 08/29/01 16:02 Rikki NMR measurements
  • 73010 08/29/01 15:59 x.jiang shift summary
  • 73005 08/29/01 14:47 apar Start_of_Run_Parity_3321
  • 73000 08/29/01 13:51 apar Start_of_Run_Parity_3320
  • 72996 08/29/01 13:48 x.jiang beam-half-plate OUT, starting run 3217, 31217.
  • 72994 08/29/01 12:56 apar Start_of_Run_Parity_3319
  • 72989 08/29/01 12:01 apar Start_of_Run_Parity_3318
  • 72985 08/29/01 11:44 x. jiang reactz of run 3217, 31217.
  • 72984 08/29/01 11:08 x. jiang spot++ run 3218,3219
  • 72982 08/29/01 11:05 apar Start_of_Run_Parity_3317
  • 72977 08/29/01 10:47 apar Start_of_Run_Parity_3316
  • 72973 08/29/01 09:51 apar Start_of_Run_Parity_3315
  • 72967 08/29/01 08:34 apar Start_of_Run_Parity_3314
  • 72962 08/29/01 08:04 A. Deur Shift summary
  • 72959 08/29/01 07:38 apar Start_of_Run_Parity_3313
  • 72953 08/29/01 06:44 apar Start_of_Run_Parity_3312
  • 72951 08/29/01 06:00 adaq Runsummary_Left_HRS
  • 72950 08/29/01 06:00 adaq Runsummary_Right_HRS
  • 72947 08/29/01 05:48 apar Start_of_Run_Parity_3311
  • 72944 08/29/01 05:19 kramerk Target - NMR/EPR - 39.7%/40.1%
  • 72942 08/29/01 04:54 apar Start_of_Run_Parity_3310
  • 72936 08/29/01 03:31 apar Start_of_Run_Parity_3309
  • 72931 08/29/01 02:36 apar Start_of_Run_Parity_3308
  • 72926 08/29/01 01:41 apar Start_of_Run_Parity_3307
  • 72923 08/29/01 00:49 A. Deur daq problem for left arm
  • 72920 08/29/01 00:45 apar Start_of_Run_Parity_3306
  • 72915 08/29/01 00:13 apar Start_of_Run_Parity_3305
  • Tuesday

  • 72912 08/28/01 23:58 S. Sirca Summary of evening shift
  • 72911 08/28/01 23:46 Xiaofeng scaler update for kin-4(Q2=1.14)
  • 72910 08/28/01 23:18 apar Start_of_Run_Parity_3304
  • 72904 08/28/01 23:04 apar Start_of_Run_Parity_3303
  • 72902 08/28/01 23:01 Xiaofeng Target NMR: 40.1% EPR: 41.3%
  • 72898 08/28/01 21:35 apar Start_of_Run_Parity_3302
  • 72892 08/28/01 20:38 apar Start_of_Run_Parity_3301
  • 72889 08/28/01 19:42 apar Start_of_Run_Parity_3300
  • 72883 08/28/01 18:47 apar Start_of_Run_Parity_3299
  • 72879 08/28/01 17:51 apar Start_of_Run_Parity_3298
  • 0 08/28/01 17:49 no_name no_key
  • 72875 08/28/01 17:40 Xiaofeng , X Target NMR: 40.9% EPR: 41.81%
  • 72873 08/28/01 16:55 apar Start_of_Run_Parity_3297
  • 72869 08/28/01 16:48 apar Start_of_Run_Parity_3296
  • 72863 08/28/01 16:01 R. Michaels day shift summary
  • 72858 08/28/01 15:23 apar Start_of_Run_Parity_3295
  • 72854 08/28/01 14:29 apar Start_of_Run_Parity_3294
  • 72853 08/28/01 14:09 R. Michaels target config change (90, wp IN)
  • 72850 08/28/01 13:26 apar Start_of_Run_Parity_3293
  • 72845 08/28/01 12:05 apar Start_of_Run_Parity_3292
  • 72842 08/28/01 11:40 R. Michaels L-arm DAQ 100% DT run 3190
  • 72839 08/28/01 11:31 apar Start_of_Run_Parity_3291
  • 72836 08/28/01 11:24 W. Korsch Target - EPR: 41.15%
  • 72833 08/28/01 11:19 apar Start_of_Run_Parity_3290
  • 72832 08/28/01 11:13 Rikki NMR measurment
  • 72828 08/28/01 09:43 apar Start_of_Run_Parity_3289
  • 72823 08/28/01 08:49 apar Start_of_Run_Parity_3288
  • 72821 08/28/01 08:38 R. Michaels OTR on target
  • 72820 08/28/01 08:03 kramerk Shift Summary
  • 72816 08/28/01 07:46 kramerk MCC needs to cycle a magnet
  • 72814 08/28/01 07:19 apar Start_of_Run_Parity_3287
  • 72809 08/28/01 06:24 apar Start_of_Run_Parity_3286
  • 72805 08/28/01 06:00 adaq Runsummary_Left_HRS
  • 72804 08/28/01 06:00 adaq Runsummary_Right_HRS
  • 72803 08/28/01 05:29 apar Start_of_Run_Parity_3285
  • 72798 08/28/01 04:53 kramerk Target - NMR/EPR - 42.2%/42.0%
  • 72795 08/28/01 04:32 apar Start_of_Run_Parity_3284
  • 72786 08/28/01 02:12 apar Start_of_Run_Parity_3282
  • 72782 08/28/01 01:20 kramerk Left Arm DAQ problems
  • 72779 08/28/01 01:17 apar Start_of_Run_Parity_3281
  • 72775 08/28/01 01:02 apar Start_of_Run_Parity_3280
  • 72772 08/28/01 00:36 Adaq Lightning impeding good beam delivery
  • Monday

  • 72771 08/27/01 23:59 S. Sirca Summary of evening shift
  • 72768 08/27/01 23:55 apar Start_of_Run_Parity_3279
  • 72763 08/27/01 23:09 apar Start_of_Run_Parity_3278
  • 72759 08/27/01 22:21 Xiaofeng Target NMR: 43.7% EPR: 43.4%
  • 72757 08/27/01 22:14 apar Start_of_Run_Parity_3277
  • 72753 08/27/01 22:07 apar Start_of_Run_Parity_3276
  • 72746 08/27/01 21:39 apar Start_of_Run_Parity_3275
  • 72742 08/27/01 20:45 apar Start_of_Run_Parity_3274
  • 72736 08/27/01 19:49 apar Start_of_Run_Parity_3273
  • 72732 08/27/01 18:55 apar Start_of_Run_Parity_3272
  • 72727 08/27/01 17:59 apar Start_of_Run_Parity_3271
  • 72722 08/27/01 17:16 Xiaofeng Target NMR: 45.6% EPR: 46.2%
  • 72720 08/27/01 17:04 apar Start_of_Run_Parity_3270
  • 72715 08/27/01 16:54 apar Start_of_Run_Parity_3269
  • 72712 08/27/01 16:03 jpchen 8/27 day shift summary
  • 72707 08/27/01 15:28 apar Start_of_Run_Parity_3268
  • 72703 08/27/01 14:34 apar Start_of_Run_Parity_3267
  • 72698 08/27/01 13:40 apar Start_of_Run_Parity_3266
  • 72695 08/27/01 13:35 Saha Arc Integral measurement
  • 72694 08/27/01 13:33 R. Michaels Y raster readback saturating
  • 72693 08/27/01 13:00 jpchen correction
  • 72692 08/27/01 12:49 jpchen Resume data taking
  • 72689 08/27/01 12:46 apar Start_of_Run_Parity_3265
  • 72688 08/27/01 12:39 atrig Trigger-Download
  • 72687 08/27/01 12:37 Olivier NMR correction
  • 72686 08/27/01 12:28 Olivier NMR measurement 43.8 %
  • 72685 08/27/01 12:28 jpchen Target: EPR=46.2%
  • 72684 08/27/01 12:25 jpchen beam spot check with OTR
  • 72683 08/27/01 12:19 R. Michaels screen shot (Gen. Tools GUI)
  • 72682 08/27/01 12:17 V. Gorbenko, Moller measurement
  • 72681 08/27/01 11:56 jpchen Beam/Moller/EPR/NMR/ Spot check
  • 72680 08/27/01 08:12 kramerk Shift Summary
  • 72678 08/27/01 06:59 kramerk MCC still having trouble with magnet
  • 72673 08/27/01 06:33 apar Start_of_Run_Parity_3264
  • 72672 08/27/01 06:27 kramerk MCC says magnet down, probably need an access
  • 72668 08/27/01 06:23 kramerk/mich Target - NMR/EPR - 44.8%/44.8%
  • 72666 08/27/01 06:20 apar Start_of_Run_Parity_3263
  • 72664 08/27/01 06:00 adaq Runsummary_Left_HRS
  • 72663 08/27/01 06:00 adaq Runsummary_Right_HRS
  • 72659 08/27/01 05:43 apar Start_of_Run_Parity_3262
  • 72656 08/27/01 04:52 kramerk Left Arm DAQ failure runs 3159 and 3160
  • 72654 08/27/01 04:49 apar Start_of_Run_Parity_3261
  • 72650 08/27/01 04:41 apar Start_of_Run_Parity_3260
  • 72647 08/27/01 04:17 kramerk Run 3158/31158 had large charge asymmetry
  • 72645 08/27/01 03:41 apar Start_of_Run_Parity_3259
  • 72641 08/27/01 03:19 apar Start_of_Run_Parity_3258
  • 72638 08/27/01 03:18 kramerk Beam Half Wave Plate moved IN
  • 72634 08/27/01 02:21 apar Start_of_Run_Parity_3257
  • 72630 08/27/01 02:08 Adaq Right HRS
  • 72628 08/27/01 01:26 apar Start_of_Run_Parity_3256
  • 72624 08/27/01 00:54 kramerk Target - NMR/EPR - 45.6%/46.0%
  • 72622 08/27/01 00:32 apar Start_of_Run_Parity_3255
  • 72619 08/27/01 00:07 F. Garibaldi Shift Summary
  • Sunday

  • 72616 08/26/01 23:20 apar Start_of_Run_Parity_3254
  • 72612 08/26/01 22:25 apar Start_of_Run_Parity_3253
  • 72607 08/26/01 21:51 Xiaofeng good electron cuts for kin-4(Q2=1.14)
  • 72605 08/26/01 21:25 apar Start_of_Run_Parity_3252
  • 72601 08/26/01 21:12 Xiaofeng BeO in left arm acceptance
  • 72599 08/26/01 20:28 apar Start_of_Run_Parity_3251
  • 72595 08/26/01 20:07 Xiaofeng Target NMR :47.1% EPR: 46.2%
  • 72593 08/26/01 19:57 apar Start_of_Run_Parity_3250
  • 72591 08/26/01 19:49 F. Garibaldi e- back to left arm
  • 72590 08/26/01 18:01 Binet Compton Tuning Summary
  • 72586 08/26/01 17:01 apar Start_of_Run_Parity_3249
  • 72584 08/26/01 16:41 averett harp scan and OTR
  • 72581 08/26/01 16:38 averett raster at 5.7 GeV
  • 72577 08/26/01 16:30 apar Start_of_Run_Parity_3248
  • 72575 08/26/01 15:55 S. Sirca Summary of day shift
  • 72574 08/26/01 14:50 Zolnierczuk target NMR=45.3/EPR=45.1%
  • 72573 08/26/01 14:23 S. Sirca Beam optics needs retune
  • 72572 08/26/01 14:17 K. McCormick Current spike
  • 72571 08/26/01 14:12 K. McCormick Current Spikes
  • 72567 08/26/01 13:39 apar Start_of_Run_Parity_3247
  • 72566 08/26/01 13:11 S. Sirca Ready for beam
  • 72564 08/26/01 10:00 S. Sirca Down-time estimate
  • 72563 08/26/01 09:14 apar Start_of_Run_Parity_3246
  • 72556 08/26/01 09:06 apar Start_of_Run_Parity_3245
  • 72555 08/26/01 09:04 S. Sirca New HRS settings
  • 72554 08/26/01 08:48 kramerk Changed disk for hbook directory to work6
  • 72552 08/26/01 08:04 j singh N2 100 psig reference cell run
  • 72551 08/26/01 07:55 J. Roche Injector trouble
  • 72550 08/26/01 07:40 A. Vacheret Shift Summary
  • 72549 08/26/01 07:39 j singh NMR/EPR Polarization Shapiro
  • 72548 08/26/01 07:18 X. Zheng Target - EPR 41.5%
  • 72547 08/26/01 06:00 adaq Runsummary_Left_HRS
  • 72546 08/26/01 06:00 adaq Runsummary_Right_HRS
  • 72542 08/26/01 05:11 apar Start_of_Run_Parity_3244
  • 72537 08/26/01 04:43 apar Start_of_Run_Parity_3243
  • 72533 08/26/01 04:32 j singh Configuration Change
  • 72531 08/26/01 04:27 apar Start_of_Run_Parity_3242
  • 72526 08/26/01 03:29 apar Start_of_Run_Parity_3241
  • 72521 08/26/01 03:00 Binet Another current spike
  • 72519 08/26/01 02:42 apar Start_of_Run_Parity_3240
  • 72514 08/26/01 01:56 apar Start_of_Run_Parity_3239
  • 72510 08/26/01 01:30 X. Zheng missing entry for EPR @ Aug-25, 8:24am
  • 72509 08/26/01 01:26 j singh Configuration Change
  • 72508 08/26/01 01:22 j singh NMR/EPR Measurement Shapiro
  • 72506 08/26/01 01:11 apar Start_of_Run_Parity_3238
  • 72504 08/26/01 01:04 X. Zheng Target - EPR 39.0%
  • 72501 08/26/01 00:07 Djawotho Swing Shift Summary
  • Saturday

  • 72499 08/25/01 23:34 apar Start_of_Run_Parity_3237
  • 72492 08/25/01 22:44 apar Start_of_Run_Parity_3236
  • 72486 08/25/01 22:00 apar Start_of_Run_Parity_3235
  • 72483 08/25/01 21:15 apar Start_of_Run_Parity_3234
  • 72479 08/25/01 20:32 Michael R, X Target NMR: 39.3%; EPR: 40.24%
  • 72477 08/25/01 20:31 apar Start_of_Run_Parity_3233
  • 72471 08/25/01 19:06 apar Start_of_Run_Parity_3232
  • 72468 08/25/01 18:59 Roedelbronn high number of good events
  • 72466 08/25/01 18:20 apar Start_of_Run_Parity_3231
  • 72460 08/25/01 17:35 apar Start_of_Run_Parity_3230
  • 72457 08/25/01 17:06 Xiaofeng Scaler update
  • 72454 08/25/01 16:51 apar Start_of_Run_Parity_3229
  • 72449 08/25/01 16:06 apar Start_of_Run_Parity_3228
  • 72446 08/25/01 16:03 S. Sirca Summary of day shift
  • 72443 08/25/01 15:22 apar Start_of_Run_Parity_3227
  • 72439 08/25/01 14:44 Zolnierczuk target EPR=40.3%/NMR=39.9%
  • 72437 08/25/01 14:36 apar Start_of_Run_Parity_3226
  • 72431 08/25/01 14:22 apar Start_of_Run_Parity_3225
  • 72430 08/25/01 14:16 Adaq previous entries refuses
  • 72429 08/25/01 13:44 S. Sirca Beam off, RF zone down
  • 72424 08/25/01 13:28 apar Start_of_Run_Parity_3224
  • 72420 08/25/01 12:42 apar Start_of_Run_Parity_3223
  • 72416 08/25/01 12:07 Zolnierczuk target 90->270 NMR=39.5%
  • 72415 08/25/01 12:03 Zolnierczuk target NMR=39.8%/EPR=41.1% (lost entry)
  • 72412 08/25/01 11:58 apar Start_of_Run_Parity_3222
  • 72410 08/25/01 11:53 apar Start_of_Run_Parity_3221
  • 72406 08/25/01 11:48 apar Start_of_Run_Parity_3220
  • 72405 08/25/01 11:36 S. Sirca Misnumbering of runs
  • 72402 08/25/01 11:17 R. Michaels problems with halog
  • 08/25/01 11:09 apar Start_of_Run_Parity_3219
  • 08/25/01 10:37 apar Start_of_Run_Parity_3218
  • 08/25/01 09:47 apar Start_of_Run_Parity_3217
  • 08/25/01 09:45 apar Start_of_Run_Parity_3216
  • 08/25/01 09:31 apar Start_of_Run_Parity_3215
  • 08/25/01 08:45 apar Start_of_Run_Parity_3214
  • 72300 08/25/01 07:58 A. Vacheret Shift Summary
  • 72299 08/25/01 07:21 apar Start_of_Run_Parity_3213
  • 72294 08/25/01 06:36 apar Start_of_Run_Parity_3212
  • 72288 08/25/01 06:00 adaq Runsummary_Left_HRS
  • 72287 08/25/01 06:00 adaq Runsummary_Right_HRS
  • 72286 08/25/01 05:49 apar Start_of_Run_Parity_3211
  • 72282 08/25/01 05:04 apar Start_of_Run_Parity_3210
  • 72277 08/25/01 04:19 apar Start_of_Run_Parity_3209
  • 72272 08/25/01 03:34 apar Start_of_Run_Parity_3208
  • 72267 08/25/01 03:04 j singh NMR and EPR Polarization for Shapiro
  • 72266 08/25/01 02:49 apar Start_of_Run_Parity_3207
  • 72263 08/25/01 02:37 X. Zheng Target - EPR 40.94%
  • 72260 08/25/01 01:25 A.Vacheret Insertable (beam) HWP OUT position.
  • 72259 08/25/01 01:17 apar Start_of_Run_Parity_3206
  • 72253 08/25/01 00:26 apar Start_of_Run_Parity_3205
  • Friday

  • 72248 08/24/01 23:55 S. Sirca Summary of evening shift
  • 72245 08/24/01 23:37 apar Start_of_Run_Parity_3204
  • 72240 08/24/01 22:52 apar Start_of_Run_Parity_3203
  • 72236 08/24/01 22:07 apar Start_of_Run_Parity_3202
  • 72231 08/24/01 22:03 apar Start_of_Run_Parity_3201
  • 72226 08/24/01 21:18 apar Start_of_Run_Parity_3200
  • 72223 08/24/01 20:48 S. Sirca Repeated CODA problems
  • 72220 08/24/01 20:34 apar Start_of_Run_Parity_3199
  • 72215 08/24/01 20:28 apar Start_of_Run_Parity_3198
  • 72211 08/24/01 20:13 Michael R, X Target NMR: 40.6% EPR: 41.1%
  • 72209 08/24/01 20:12 apar Start_of_Run_Parity_3197
  • 72204 08/24/01 19:26 apar Start_of_Run_Parity_3196
  • 72199 08/24/01 18:41 apar Start_of_Run_Parity_3195
  • 72194 08/24/01 17:56 apar Start_of_Run_Parity_3194
  • 72191 08/24/01 17:52 S. Sirca ROC14, ROC2 reboots
  • 72188 08/24/01 17:45 apar Start_of_Run_Parity_3193
  • 72185 08/24/01 17:37 apar Start_of_Run_Parity_3192
  • 72182 08/24/01 17:12 Roedelbronn Left HRS
  • 72180 08/24/01 16:42 apar Start_of_Run_Parity_3191
  • 72175 08/24/01 16:05 Higinbotham Shift Summary
  • 72172 08/24/01 15:56 apar Start_of_Run_Parity_3190
  • 72168 08/24/01 15:12 apar Start_of_Run_Parity_3189
  • 72163 08/24/01 14:21 Zolnierczuk target NMR=41.5%/EPR=42.9%
  • 72162 08/24/01 12:53 Higinbotham He Levels OK, Restarting Magnets
  • 72161 08/24/01 10:58 Zolnierczuk target NMR=40.2%/EPR=41.2%
  • 72160 08/24/01 10:41 Higinbotham ESR Trip
  • 72156 08/24/01 10:05 apar Start_of_Run_Parity_3187
  • 72147 08/24/01 08:34 apar Start_of_Run_Parity_3185
  • 72142 08/24/01 08:02 A. Vacheret Shift summary
  • 72139 08/24/01 07:44 apar Start_of_Run_Parity_3184
  • 72135 08/24/01 07:20 kramerk Target - NMR/EPR - 40.6%/41.3%
  • 72132 08/24/01 07:00 apar Start_of_Run_Parity_3183
  • 72131 08/24/01 06:52 Olivier NMR measurement
  • 72128 08/24/01 06:21 A. Vacheret Charge asymmetry
  • 72127 08/24/01 06:00 adaq Runsummary_Left_HRS
  • 72126 08/24/01 06:00 adaq Runsummary_Right_HRS
  • 72123 08/24/01 05:43 apar Start_of_Run_Parity_3182
  • 72118 08/24/01 04:58 apar Start_of_Run_Parity_3181
  • 72115 08/24/01 04:16 Olivier Holding filed change
  • 72113 08/24/01 04:14 apar Start_of_Run_Parity_3180
  • 72108 08/24/01 03:13 apar Start_of_Run_Parity_3179
  • 72103 08/24/01 02:27 apar Start_of_Run_Parity_3178
  • 72099 08/24/01 01:31 apar Start_of_Run_Parity_3177
  • 72095 08/24/01 00:46 kramerk/gayo Target - NMR/EPR - 41.0%/42.1%
  • 72094 08/24/01 00:40 apar Start_of_Run_Parity_3176
  • 72088 08/24/01 00:01 slifer shift summary
  • Thursday

  • 72085 08/23/01 23:19 apar Start_of_Run_Parity_3175
  • 72080 08/23/01 22:35 apar Start_of_Run_Parity_3174
  • 72074 08/23/01 21:51 apar Start_of_Run_Parity_3173
  • 72070 08/23/01 21:07 apar Start_of_Run_Parity_3172
  • 72065 08/23/01 20:22 apar Start_of_Run_Parity_3171
  • 72060 08/23/01 19:38 apar Start_of_Run_Parity_3170
  • 72057 08/23/01 18:55 Pibero, Xiao Target NMR: 41.7%, EPR: 42.9%
  • 72055 08/23/01 18:53 apar Start_of_Run_Parity_3169
  • 72050 08/23/01 17:27 apar Start_of_Run_Parity_3168
  • 72045 08/23/01 16:43 apar Start_of_Run_Parity_3167
  • 72041 08/23/01 16:35 S. Sirca LN2 level for HRS left
  • 72039 08/23/01 15:59 Higinbotham Shift Summary
  • 72037 08/23/01 15:58 apar Start_of_Run_Parity_3166
  • 72033 08/23/01 15:56 Zolnierczuk Gas Cerenkov update
  • 72031 08/23/01 15:13 apar Start_of_Run_Parity_3165
  • 72026 08/23/01 14:26 apar Start_of_Run_Parity_3164
  • 72022 08/23/01 14:05 averett BeO in left acceptance
  • 72021 08/23/01 13:58 Zolnierczuk target rotated to 270 degree
  • 72020 08/23/01 13:55 Higinbotham Beam Half Wave In Starting 3054
  • 72019 08/23/01 13:42 apar Start_of_Run_Parity_3163
  • 72014 08/23/01 13:22 apar Start_of_Run_Parity_3162
  • 72008 08/23/01 13:15 apar Start_of_Run_Parity_3161
  • 72004 08/23/01 13:09 Zolnierczuk target RTD#5
  • 72003 08/23/01 13:06 apar Start_of_Run_Parity_3160
  • 71998 08/23/01 12:52 Zolnierczuk Gas Cerenkov calibration
  • 71997 08/23/01 12:11 apar Start_of_Run_Parity_3159
  • 71992 08/23/01 11:42 Higinbotham Magnet Screen (Kinematics#3)
  • 71991 08/23/01 11:36 Zolnierczuk target EPR=43.3/NMR=42.7%
  • 71989 08/23/01 11:25 apar Start_of_Run_Parity_3158
  • 71984 08/23/01 10:51 apar Start_of_Run_Parity_3157
  • 71981 08/23/01 10:30 apar Start_of_Run_Parity_3156
  • 71973 08/23/01 09:41 apar Start_of_Run_Parity_3155
  • 71970 08/23/01 09:10 Xiaofeng Target EPR:43.5%
  • 71969 08/23/01 08:56 apar Start_of_Run_Parity_3154
  • 71964 08/23/01 08:17 A. Vacheret shift summary
  • 71963 08/23/01 08:14 A. vacheret charge asymmetry summary plot
  • 71961 08/23/01 08:07 apar Start_of_Run_Parity_3153
  • 71956 08/23/01 07:52 Xiaofeng good electron cuts for kin-3
  • 71953 08/23/01 07:11 apar Start_of_Run_Parity_3152
  • 71949 08/23/01 06:26 apar Start_of_Run_Parity_3151
  • 71945 08/23/01 06:00 adaq Runsummary_Left_HRS
  • 71944 08/23/01 06:00 adaq Runsummary_Right_HRS
  • 71943 08/23/01 05:53 J. Roche Spectrometers setting for kinematic 3
  • 71940 08/23/01 05:42 apar Start_of_Run_Parity_3150
  • 71939 08/23/01 05:41 Xiaofeng For kin-3, run starts at 3050/31050
  • 71937 08/23/01 05:38 Olivier kinematic 2 summary
  • 71936 08/23/01 05:27 Julie, Xiaof Target NMR :42.8% EPR : 43.0%
  • 71935 08/23/01 05:21 A. Vacheret Spectrometers motion control problem fixed
  • 71934 08/23/01 04:48 A. Vacheret Spectrometers motion GUI doesn't work
  • 71933 08/23/01 04:16 Adaq left arm bogie error message
  • 71930 08/23/01 02:44 apar Start_of_Run_Parity_3149
  • 71925 08/23/01 01:56 apar Start_of_Run_Parity_3149
  • 71923 08/23/01 01:53 apar Start_of_Run_Parity_3148
  • 71918 08/23/01 01:07 apar Start_of_Run_Parity_3147
  • 71914 08/23/01 00:31 jpchen 8/22 swing shift summary
  • 71913 08/23/01 00:31 jpchen target: EPR=42.9%
  • 71911 08/23/01 00:22 apar Start_of_Run_Parity_3146
  • 71909 08/23/01 00:20 Djawotho/Che Target-NMR=41.92%
  • Wednesday

  • 71905 08/22/01 23:24 jpchen beam current fluctation
  • 71902 08/22/01 23:05 apar Start_of_Run_Parity_3145
  • 71898 08/22/01 22:21 apar Start_of_Run_Parity_3144
  • 71892 08/22/01 21:37 apar Start_of_Run_Parity_3143
  • 71890 08/22/01 21:12 Binet Update on Compton polarimeter
  • 71888 08/22/01 20:53 apar Start_of_Run_Parity_3142
  • 71883 08/22/01 20:09 apar Start_of_Run_Parity_3141
  • 71877 08/22/01 19:24 apar Start_of_Run_Parity_3140
  • 71876 08/22/01 19:06 jpchen/piber target: EPR=44.7%
  • 71875 08/22/01 19:03 Djawotho Target-NMR=44.22%
  • 71872 08/22/01 18:53 jpchen beam back, resume data taking
  • 71871 08/22/01 18:53 apar Start_of_Run_Parity_3139
  • 71868 08/22/01 17:00 Higinbotham Shift Summary
  • 71866 08/22/01 15:05 Zolnierczuk TARGET EPR=42.9%/NMR=41.6%
  • 71863 08/22/01 13:43 apar Start_of_Run_Parity_3138
  • 71862 08/22/01 13:41 Zolnierczuk reference cell He-3 loss update
  • 71857 08/22/01 12:35 apar Start_of_Run_Parity_3137
  • 71852 08/22/01 11:51 Zolnierczuk target: reference cell He-3 pressure loss
  • 71851 08/22/01 11:49 apar Start_of_Run_Parity_3136
  • 71846 08/22/01 11:08 Higinbotham Right GUI Fixed !?
  • 71842 08/22/01 10:59 Higinbotham Magnet Screen
  • 71841 08/22/01 10:33 Binet Monitoring of the beam polarization
  • 71837 08/22/01 09:55 apar Start_of_Run_Parity_3134
  • 71835 08/22/01 09:07 Zolnierczuk target EPR=42.2%/NMR=40.9%
  • 71825 08/22/01 08:01 kramerk Shift Summary
  • 71822 08/22/01 07:21 apar Start_of_Run_Parity_3131
  • 71818 08/22/01 06:36 apar Start_of_Run_Parity_3130
  • 71813 08/22/01 06:00 adaq Runsummary_Left_HRS
  • 71812 08/22/01 06:00 adaq Runsummary_Right_HRS
  • 71811 08/22/01 05:51 apar Start_of_Run_Parity_3129
  • 71805 08/22/01 05:06 apar Start_of_Run_Parity_3128
  • 71801 08/22/01 04:21 apar Start_of_Run_Parity_3127
  • 71795 08/22/01 03:36 apar Start_of_Run_Parity_3126
  • 71791 08/22/01 02:51 apar Start_of_Run_Parity_3125
  • 71786 08/22/01 02:29 kramerk Target - NMR/EPR - 41.3%/42.5%
  • 71780 08/22/01 00:51 apar Start_of_Run_Parity_3123
  • 71776 08/22/01 00:18 kramerk Target - Laser #4 not turned on
  • 71771 08/22/01 00:01 Reitz Shift summary
  • Tuesday

  • 71762 08/21/01 22:36 apar Start_of_Run_Parity_3120
  • 71759 08/21/01 22:09 Pibero , Xia Target-NMR: 43.4%, EPR: 43.6%
  • 71758 08/21/01 21:50 apar Start_of_Run_Parity_3119
  • 71751 08/21/01 20:27 apar Start_of_Run_Parity_3118
  • 71748 08/21/01 20:13 Reitz LN_2 Level left dipole
  • 71747 08/21/01 19:42 apar Start_of_Run_Parity_3117
  • 71741 08/21/01 18:58 apar Start_of_Run_Parity_3116
  • 71737 08/21/01 18:40 Reitz Beam off for 10 min
  • 71736 08/21/01 18:04 Higinbotham Right GUI Not Working Again
  • 71735 08/21/01 17:58 apar Start_of_Run_Parity_3115
  • 71730 08/21/01 17:12 Pibero/Xiaof Target-NMR=44.5%,EPR=44.8%
  • 71727 08/21/01 17:09 apar Start_of_Run_Parity_3114
  • 71726 08/21/01 17:06 Reitz magnets on right arm are fine
  • 71723 08/21/01 16:58 apar Start_of_Run_Parity_3113
  • 71720 08/21/01 16:49 reitz Start and End of Run Scripts fail
  • 71719 08/21/01 16:45 Bodo Reitz DAQ problems
  • 71718 08/21/01 16:43 apar Start_of_Run_Parity_3112
  • 71715 08/21/01 16:42 apar Start_of_Run_Parity_3111
  • 71706 08/21/01 16:12 Higinbotham Shift Summary
  • 71704 08/21/01 16:06 Higinbotham Magnet Trips
  • 71703 08/21/01 13:17 apar Start_of_Run_Parity_3110
  • 71698 08/21/01 12:16 apar Start_of_Run_Parity_2109
  • 71695 08/21/01 12:14 Higinbotham Right Arm GUI Fixed
  • 71692 08/21/01 12:02 apar Start_of_Run_Parity_2108
  • 71689 08/21/01 11:54 apar Start_of_Run_Parity_2107
  • 71670 08/21/01 11:12 apar Start_of_Run_Parity_2106
  • 71665 08/21/01 11:00 Zolnierczuk target EPR=41.7%/NMR=40.8%
  • 71663 08/21/01 09:24 apar Start_of_Run_Parity_2104
  • 71658 08/21/01 08:39 apar Start_of_Run_Parity_2103
  • 71651 08/21/01 08:32 apar Start_of_Run_Parity_2102
  • 71648 08/21/01 08:15 A .Vacheret shift summary
  • 71641 08/21/01 07:02 apar Start_of_Run_Parity_2100
  • 71636 08/21/01 06:16 apar Start_of_Run_Parity_2099
  • 71631 08/21/01 06:00 adaq Runsummary_Left_HRS
  • 71630 08/21/01 06:00 adaq Runsummary_Right_HRS
  • 71627 08/21/01 05:31 apar Start_of_Run_Parity_2098
  • 71624 08/21/01 04:51 kramerk Target - NMR Polarization = 41.4%
  • 71623 08/21/01 04:45 X. Zheng Target - EPR 42.25%
  • 71621 08/21/01 04:41 apar Start_of_Run_Parity_2097
  • 71619 08/21/01 04:32 kramerk Changed hbook link to /adaql1/work5
  • 71614 08/21/01 03:28 apar Start_of_Run_Parity_2096
  • 71610 08/21/01 02:42 apar Start_of_Run_Parity_2095
  • 71605 08/21/01 01:57 apar Start_of_Run_Parity_2094
  • 71600 08/21/01 01:12 apar Start_of_Run_Parity_2093
  • 71594 08/21/01 00:27 apar Start_of_Run_Parity_2092
  • 71591 08/21/01 00:19 A. Deur Shift summary
  • Monday

  • 71587 08/20/01 23:43 apar Start_of_Run_Parity_2091
  • 71582 08/20/01 22:58 apar Start_of_Run_Parity_2090
  • 71578 08/20/01 22:31 A. Tobias Target - NMR: 41.97%
  • 71577 08/20/01 22:26 W. Korsch Target - EPR: 42.5%
  • 71575 08/20/01 22:13 apar Start_of_Run_Parity_2089
  • 71571 08/20/01 21:03 A. Deur Right Arm HV
  • 71568 08/20/01 20:56 apar Start_of_Run_Parity_2088
  • 71564 08/20/01 20:40 apar Start_of_Run_Parity_2087
  • 71561 08/20/01 19:58 A. Deur roc14
  • 71559 08/20/01 19:55 apar Start_of_Run_Parity_2086
  • 71556 08/20/01 19:23 A. Deur bcm problem for runs 2(1)973
  • 71554 08/20/01 19:09 apar Start_of_Run_Parity_2085
  • 71544 08/20/01 17:42 apar Start_of_Run_Parity_2083
  • 71539 08/20/01 16:57 apar Start_of_Run_Parity_2082
  • 71535 08/20/01 16:21 Higinbotham ARC Energy (BDL only)
  • 71533 08/20/01 16:19 Higinbotham Shift Summary
  • 71530 08/20/01 16:11 apar Start_of_Run_Parity_2081
  • 71527 08/20/01 15:49 jpchen target:NMR=43.8%, EPR=44.3%
  • 71526 08/20/01 15:25 apar Start_of_Run_Parity_2080
  • 71523 08/20/01 15:08 gilman target polarization orientation correction
  • 71520 08/20/01 14:15 apar Start_of_Run_Parity_2079
  • 71514 08/20/01 13:22 apar Start_of_Run_Parity_2078
  • 71511 08/20/01 13:04 R. Michaels PZT test finished
  • 71510 08/20/01 13:03 R. Suleiman PZT test
  • 71509 08/20/01 12:55 apar Start_of_Run_Parity_2077
  • 71508 08/20/01 12:43 apar Start_of_Run_Parity_2076
  • 71506 08/20/01 12:16 apar Start_of_Run_Parity_2075
  • 71501 08/20/01 11:52 apar Start_of_Run_Parity_2074
  • 71498 08/20/01 11:35 R. Michaels PZT mirror tests
  • 71495 08/20/01 11:14 apar Start_of_Run_Parity_2073
  • 71487 08/20/01 10:53 apar Start_of_Run_Parity_2072
  • 71482 08/20/01 10:47 apar Start_of_Run_Parity_2071
  • 71479 08/20/01 10:44 gilman magnets back
  • 71478 08/20/01 10:24 jpchen Target: NMR=41.4%, EPR=42.0%
  • 71477 08/20/01 09:59 gilman right dipole problem, quad 3
  • 71476 08/20/01 09:21 gilman Quadrupole trips
  • 71473 08/20/01 08:19 apar Start_of_Run_Parity_2070
  • 71469 08/20/01 08:10 kramerk Shift Summary
  • 71466 08/20/01 07:34 apar Start_of_Run_Parity_2069
  • 71460 08/20/01 06:49 apar Start_of_Run_Parity_2068
  • 71455 08/20/01 06:05 apar Start_of_Run_Parity_2067
  • 71452 08/20/01 06:00 adaq Runsummary_Left_HRS
  • 71451 08/20/01 06:00 adaq Runsummary_Right_HRS
  • 71448 08/20/01 05:20 apar Start_of_Run_Parity_2066
  • 71444 08/20/01 05:15 kramerk Another Current spike!
  • 71442 08/20/01 04:35 apar Start_of_Run_Parity_2065
  • 71439 08/20/01 04:01 kramerk/mcco Target - NMR/EPR - 42.2%/42.2%
  • 71436 08/20/01 03:50 apar Start_of_Run_Parity_2064
  • 71431 08/20/01 02:35 apar Start_of_Run_Parity_2063
  • 71427 08/20/01 01:51 apar Start_of_Run_Parity_2062
  • 71423 08/20/01 01:05 kramerk Beam is back!!
  • 71422 08/20/01 01:05 apar Start_of_Run_Parity_2061
  • 71418 08/20/01 00:21 A. Deur Shift summary
  • 71416 08/20/01 00:01 A. Deur Beam down.
  • Sunday

  • 71414 08/19/01 23:07 R. Michaels R-arm DAQ fix, no TCL dialogs
  • 71411 08/19/01 22:57 apar Start_of_Run_Parity_2060
  • 71405 08/19/01 22:37 apar Start_of_Run_Parity_2059
  • 71404 08/19/01 22:30 A. Deur coda problem
  • 71401 08/19/01 22:24 A. Deur eP energy measurement
  • 71400 08/19/01 22:16 Xiaofeng right arm preshower/shower calibrations
  • 71396 08/19/01 22:09 apar Start_of_Run_Parity_2058
  • 71392 08/19/01 22:04 apar Start_of_Run_Parity_2057
  • 71388 08/19/01 22:01 apar Start_of_Run_Parity_2056
  • 71386 08/19/01 21:50 A. Tobias, X Target - NMR: 42.53%, EPR: 42.8%
  • 71385 08/19/01 21:47 apar Start_of_Run_Parity_2055
  • 71374 08/19/01 20:44 A. Tobias Target - NMR: 41.08%
  • 71368 08/19/01 20:21 A. Deur beam is back
  • 71367 08/19/01 20:03 A. Deur Beam down
  • 71366 08/19/01 19:40 W. Korsch Target - EPR: 41.25%
  • 71364 08/19/01 18:23 apar Start_of_Run_Parity_2054
  • 71358 08/19/01 17:42 A. Deur ROC14
  • 71355 08/19/01 17:37 apar Start_of_Run_Parity_2053
  • 71353 08/19/01 16:47 apar Start_of_Run_Parity_2052
  • 71349 08/19/01 16:05 Ransome shift summary - day shift
  • 71347 08/19/01 16:01 apar Start_of_Run_Parity_2051
  • 71342 08/19/01 15:14 apar Start_of_Run_Parity_2050
  • 71335 08/19/01 14:54 apar Start_of_Run_Parity_2049
  • 71330 08/19/01 14:25 apar Start_of_Run_Parity_2048
  • 71326 08/19/01 13:42 Xiaofeng, Pi Target NMR : 41.6% EPR: 41.8%
  • 71319 08/19/01 12:17 apar Start_of_Run_Parity_2046
  • 71314 08/19/01 11:30 apar Start_of_Run_Parity_2045
  • 71310 08/19/01 10:45 apar Start_of_Run_Parity_2044
  • 71304 08/19/01 09:59 apar Start_of_Run_Parity_2043
  • 71299 08/19/01 09:13 apar Start_of_Run_Parity_2042
  • 71295 08/19/01 08:40 X. Zheng Target rotate to 90 degree (late entry)
  • 71294 08/19/01 08:38 X. Zheng Target - NMR 42.02%, EPR 41.82%
  • 71292 08/19/01 08:27 apar Start_of_Run_Parity_2041
  • 71289 08/19/01 08:03 W. Korsch Owl shift summary
  • 71284 08/19/01 07:15 apar Start_of_Run_Parity_2040
  • 71281 08/19/01 06:52 K. McCormick Another Beam Current Spike
  • 71278 08/19/01 06:28 apar Start_of_Run_Parity_2039
  • 71275 08/19/01 06:00 adaq Runsummary_Left_HRS
  • 71274 08/19/01 06:00 adaq Runsummary_Right_HRS
  • 71273 08/19/01 05:47 W. Korsch Target field at 90 degrees
  • 71270 08/19/01 05:43 apar Start_of_Run_Parity_2038
  • 71265 08/19/01 04:59 apar Start_of_Run_Parity_2037
  • 71260 08/19/01 04:14 apar Start_of_Run_Parity_2036
  • 71257 08/19/01 03:24 W. Korsch Beam current spike
  • 71254 08/19/01 03:12 apar Start_of_Run_Parity_2035
  • 71249 08/19/01 02:27 apar Start_of_Run_Parity_2034
  • 71246 08/19/01 02:00 X. Zheng Target - NMR 42.81%, EPR 43.3%
  • 71243 08/19/01 01:43 apar Start_of_Run_Parity_2033
  • 71242 08/19/01 01:17 K. McCormick Changed max number of events for parity run
  • 71237 08/19/01 00:30 apar Start_of_Run_Parity_2032
  • 71234 08/19/01 00:08 A. Deur shift summary
  • Saturday

  • 71231 08/18/01 23:49 apar Start_of_Run_Parity_2031
  • 71227 08/18/01 23:13 A. Deur Elastic cross sections
  • 71225 08/18/01 23:08 apar Start_of_Run_Parity_2030
  • 71220 08/18/01 22:26 apar Start_of_Run_Parity_2029
  • 71215 08/18/01 21:44 apar Start_of_Run_Parity_2028
  • 71210 08/18/01 21:03 apar Start_of_Run_Parity_2027
  • 71208 08/18/01 19:39 kramerk,tobi Target -NMR/EPR - 43.2%/43.0%
  • 71204 08/18/01 19:11 apar Start_of_Run_Parity_2026
  • 71200 08/18/01 17:55 apar Start_of_Run_Parity_2025
  • 71194 08/18/01 17:14 apar Start_of_Run_Parity_2024
  • 71189 08/18/01 16:31 apar Start_of_Run_Parity_2023
  • 71184 08/18/01 16:02 ransome shift summary
  • 71183 08/18/01 15:58 Binet Vertical Scan in the Compton Chicane (late entry)
  • 71181 08/18/01 15:48 apar Start_of_Run_Parity_2022
  • 71177 08/18/01 15:28 ransome good event decline, parity trigger problem?
  • 71175 08/18/01 15:06 apar Start_of_Run_Parity_2021
  • 71170 08/18/01 14:25 apar Start_of_Run_Parity_2020
  • 71165 08/18/01 13:43 apar Start_of_Run_Parity_2019
  • 71161 08/18/01 13:24 Xiaofeng, Pi target NMR:43.99% EPR: 43.7%
  • 71159 08/18/01 13:02 apar Start_of_Run_Parity_2018
  • 71154 08/18/01 11:42 apar Start_of_Run_Parity_2017
  • 71149 08/18/01 11:00 apar Start_of_Run_Parity_2016
  • 71144 08/18/01 10:19 apar Start_of_Run_Parity_2015
  • 71139 08/18/01 09:38 apar Start_of_Run_Parity_2014
  • 71135 08/18/01 09:21 Djawotho NMR & EPR
  • 71134 08/18/01 09:10 X. Zheng Target - EPR: 45.1%
  • 71132 08/18/01 08:56 apar Start_of_Run_Parity_2013
  • 71128 08/18/01 08:04 W. Korsch Owl shift summary
  • 71124 08/18/01 07:46 apar Start_of_Run_Parity_2012
  • 71119 08/18/01 07:32 apar Start_of_Run_Parity_2011
  • 71114 08/18/01 06:50 apar Start_of_Run_Parity_2010
  • 71109 08/18/01 06:09 apar Start_of_Run_Parity_2009
  • 71108 08/18/01 06:08 W. Korsch Beam back
  • 71107 08/18/01 06:00 adaq Runsummary_Left_HRS
  • 71106 08/18/01 06:00 adaq Runsummary_Right_HRS
  • 71103 08/18/01 04:46 W. Korsch Problems with beam due to thunderstorm
  • 71100 08/18/01 04:30 apar Start_of_Run_Parity_2008
  • 71095 08/18/01 03:49 apar Start_of_Run_Parity_2007
  • 71094 08/18/01 03:49 W. Korsch More RF problems
  • 71089 08/18/01 02:56 apar Start_of_Run_Parity_2006
  • 71088 08/18/01 02:47 X. Zheng Target - NMR 47.02%, EPR 46.03%
  • 71087 08/18/01 02:44 W. Korsch RF fault
  • 71086 08/18/01 02:14 W. Korsch Q2L works again
  • 71085 08/18/01 00:29 W. Korsch Q2L quenched
  • 71080 08/18/01 00:02 A. Deur Beam Summary
  • Friday

  • 71073 08/17/01 23:17 apar Start_of_Run_Parity_2004
  • 71064 08/17/01 22:12 Eugene Moller results
  • 71063 08/17/01 21:54 Xiaofeng, Da Target NMR:48.0% EPR: 46.9%
  • 71061 08/17/01 21:53 apar Start_of_Run_Parity_2002
  • 71057 08/17/01 21:36 apar Start_of_Run_Parity_2001
  • 71054 08/17/01 21:29 Armstrong NMR -> 48.0%
  • 71053 08/17/01 17:53 Xiaofeng, Da Target NMR: 46.9% EPR: 45.8%
  • 71048 08/17/01 16:44 apar Start_of_Run_Parity_2000
  • 71045 08/17/01 16:24 Armstrong MCC Elog entry on current fluctuations
  • 71044 08/17/01 16:10 Armstrong Beam is back!
  • 71042 08/17/01 16:04 gilman shift summary - Aug 17 day
  • 71040 08/17/01 16:03 apar Start_of_Run_Parity_1999
  • 71031 08/17/01 15:06 Sulkosky target- Holding field now 270 degrees.
  • 71030 08/17/01 15:05 Binet Update on the Compton steering
  • 71027 08/17/01 15:05 apar Start_of_Run_Parity_1998
  • 71026 08/17/01 15:03 Xiaofeng, Target NMR: 48.2% EPR: 47.2%
  • 71021 08/17/01 14:30 apar Start_of_Run_Parity_1997
  • 71017 08/17/01 14:02 apar Start_of_Run_Parity_1996
  • 71011 08/17/01 13:22 apar Start_of_Run_Parity_1995
  • 71003 08/17/01 12:14 apar Start_of_Run_Parity_1994
  • 70998 08/17/01 11:33 apar Start_of_Run_Parity_1993
  • 70995 08/17/01 11:07 Nilanga Change in ADC peak position
  • 70992 08/17/01 11:01 apar Start_of_Run_Parity_1992
  • 70991 08/17/01 10:50 finn Moller run scheduled for 5 pm
  • 70988 08/17/01 10:42 apar Start_of_Run_Parity_1991
  • 70983 08/17/01 10:27 sulkosky Target- rotate back to 0 degrees.
  • 70982 08/17/01 10:03 Sulkosky Target- Configuration change
  • 70981 08/17/01 10:01 Wolfgang/Sul Target - NMR: EPR: 45.7%
  • 70978 08/17/01 10:00 apar Start_of_Run_Parity_1990
  • 70973 08/17/01 09:20 apar Start_of_Run_Parity_1989
  • 70967 08/17/01 08:46 apar Start_of_Run_Parity_1988
  • 70964 08/17/01 08:34 Binet Compton chicane steering
  • 70963 08/17/01 08:13 Binet Compton chicane steering
  • 70960 08/17/01 08:06 apar Start_of_Run_Parity_1987
  • 70958 08/17/01 08:04 J. Roche shift summary
  • 70956 08/17/01 08:00 Binet Polarimetry measurement at E=3469 MeV
  • 70954 08/17/01 07:26 apar Start_of_Run_Parity_1986
  • 70950 08/17/01 06:57 kramerk Target - NMR/EPR - 46.9%/46.9%
  • 70948 08/17/01 06:44 apar Start_of_Run_Parity_1985
  • 70943 08/17/01 06:00 adaq Runsummary_Left_HRS
  • 70942 08/17/01 06:00 adaq Runsummary_Right_HRS
  • 70941 08/17/01 05:57 J. Roche bad beam position for run 2863/21863
  • 70939 08/17/01 05:50 apar Start_of_Run_Parity_1984
  • 70934 08/17/01 05:08 apar Start_of_Run_Parity_1983
  • 70929 08/17/01 04:27 apar Start_of_Run_Parity_1982
  • 70924 08/17/01 03:46 apar Start_of_Run_Parity_1981
  • 70920 08/17/01 03:05 apar Start_of_Run_Parity_1980
  • 70915 08/17/01 02:13 apar Start_of_Run_Parity_1979
  • 70911 08/17/01 01:57 kramerk Target - NMR/EPR - 47.5/46.5%
  • 70909 08/17/01 01:31 apar Start_of_Run_Parity_1978
  • 70907 08/17/01 01:18 Antonin charge asym slope calibration
  • 70902 08/17/01 01:04 apar Start_of_Run_Parity_1977
  • 70897 08/17/01 00:59 apar Start_of_Run_Parity_1976
  • 70892 08/17/01 00:41 apar Start_of_Run_Parity_1975
  • 70889 08/17/01 00:30 A. Deur Shift summary
  • 70886 08/17/01 00:26 apar Start_of_Run_Parity_1974
  • 70880 08/17/01 00:15 apar Start_of_Run_Parity_1973
  • 70875 08/17/01 00:03 apar Start_of_Run_Parity_1972
  • Thursday

  • 70870 08/16/01 23:53 apar Start_of_Run_Parity_1971
  • 70865 08/16/01 23:44 apar Start_of_Run_Parity_1970
  • 70861 08/16/01 23:18 A. Deur half wave in
  • 70860 08/16/01 23:18 apar Start_of_Run_Parity_1969
  • 70854 08/16/01 22:34 apar Start_of_Run_Parity_1968
  • 70849 08/16/01 21:52 apar Start_of_Run_Parity_1967
  • 70844 08/16/01 21:12 apar Start_of_Run_Parity_1966
  • 70841 08/16/01 20:13 A. Deur Beam down because of BLM alarm
  • 70840 08/16/01 20:11 A. Deur ADC peak position
  • 70839 08/16/01 19:51 A. Deur Shower/preshower
  • 70835 08/16/01 19:17 A. Deur FSD
  • 70834 08/16/01 19:11 busato end of parity run #1964 failed
  • 70828 08/16/01 18:52 A. Deur beam down for Hall B tuning
  • 70827 08/16/01 18:44 Xiaofeng ,Co Target NMR:48.5% EPR: 47.3%
  • 70824 08/16/01 18:38 apar Start_of_Run_Parity_1964
  • 70823 08/16/01 18:35 A. Deur y target distribution
  • 70820 08/16/01 17:58 A. Deur y target distribution
  • 70819 08/16/01 17:44 X. Zheng Gas Cherenkov Calibration for E99117/E97103
  • 70816 08/16/01 17:29 apar Start_of_Run_Parity_1963
  • 70813 08/16/01 16:49 A. Deur Beam tunning
  • 70810 08/16/01 16:48 apar Start_of_Run_Parity_1962
  • 70805 08/16/01 16:04 apar Start_of_Run_Parity_1961
  • 70801 08/16/01 15:33 averett Day Shift Summary
  • 70798 08/16/01 15:24 apar Start_of_Run_Parity_1960
  • 70795 08/16/01 15:11 Armstrong IOC reboot affecting BPMs
  • 70792 08/16/01 14:48 apar Start_of_Run_Parity_1959
  • 70789 08/16/01 14:11 averett beam current oscillation
  • 70786 08/16/01 14:03 apar Start_of_Run_Parity_1958
  • 70783 08/16/01 13:41 averett beam unstable
  • 70782 08/16/01 13:26 averett production running
  • 70779 08/16/01 13:23 apar Start_of_Run_Parity_1957
  • 70778 08/16/01 12:43 Armstrong NMR - 51.0%
  • 70775 08/16/01 12:35 armstrong raster setup for 4.6 GeV
  • 70774 08/16/01 12:25 averett OTR and Harp for 4.6 GeV
  • 70771 08/16/01 12:15 apar Start_of_Run_Parity_1956
  • 70766 08/16/01 11:14 apar Start_of_Run_Parity_1955
  • 70765 08/16/01 10:56 averett pulsed beam
  • 70764 08/16/01 09:34 X. Zheng Target - EPR 48.5%
  • 70763 08/16/01 09:20 Armstrong NMR - 50.3% :-)
  • 70762 08/16/01 08:07 W. Korsch Owl shift summary
  • 70761 08/16/01 07:14 W. Korsch Beam in Hall
  • 70760 08/16/01 07:02 Roedelbronn Tuning beam
  • 70759 08/16/01 06:00 adaq Runsummary_Left_HRS
  • 70758 08/16/01 06:00 adaq Runsummary_Right_HRS
  • 70757 08/16/01 04:25 W. Korsch Spectrometers set to 1.99 GeV
  • 70756 08/16/01 03:55 W. Korsch Update on Q2L
  • 70755 08/16/01 02:16 W. Korsch Update on magnet quenches
  • 70754 08/16/01 01:57 Xiaofeng, Wo Target NMR: 48.5% EPR: 47.3%
  • 70753 08/16/01 01:08 Roedelbronn NMR measurement 48.5%
  • 70752 08/16/01 01:07 jpchen magnets/ESR update
  • Wednesday

  • 70749 08/15/01 23:11 jpchen Hall A in beam permit since 10 pm
  • 70748 08/15/01 23:06 jpchen Helium clow very low, all magnets tripped off
  • 70747 08/15/01 21:16 jpchen Transverse window (on the scattering chamber) replaced
  • 70746 08/15/01 19:52 X. Zheng Target - More EPRs, latest 48.7%
  • 70745 08/15/01 18:30 jpchen Target polarization: NMR=49.9%, EPR=50.1%
  • 70744 08/15/01 17:12 jpchen Window on the scattering chamber for transver lasers
  • 70743 08/15/01 17:08 jpchen Hall A sweeped and in controlled access now
  • 70741 08/15/01 15:48 apar Start_of_Run_Parity_1954
  • 70738 08/15/01 15:34 apar Start_of_Run_Parity_1953
  • 70735 08/15/01 15:18 averett laser spectrum now updating
  • 70734 08/15/01 15:11 averett hall closed
  • 70733 08/15/01 14:58 averett Hall A target PC
  • 70732 08/15/01 10:16 W. Korsch Reference Cell pumping
  • 70731 08/15/01 10:15 W. Korsch Target - EPR: 45.5%
  • 70730 08/15/01 10:09 Armstrong NMR - 46.7% :-)
  • 70729 08/15/01 06:00 adaq Runsummary_Left_HRS
  • 70728 08/15/01 06:00 adaq Runsummary_Right_HRS
  • Tuesday

  • 70727 08/14/01 18:35 jpchen Laser wave length checked and adjusted
  • 70726 08/14/01 16:57 wimbish Gas Cerenkov CO2 SFC grade replaced with Coleman grade
  • 70725 08/14/01 16:16 Roedelbronn laser temperatures
  • 70724 08/14/01 11:31 Armstrong left dipole N2 level
  • 70723 08/14/01 09:44 Armstrong cooling jets flow on
  • 70722 08/14/01 09:43 W. Korsch Target - EPR: 38.7%
  • 70721 08/14/01 09:42 Armstrong NMR 35.4% (lasers on)
  • 70720 08/14/01 07:08 kramerk Preliminary right arm electron asymmetries for Q2=0.58
  • 70719 08/14/01 07:07 kramerk Preliminary left arm electron asymmetries for Q2=0.58
  • 70718 08/14/01 06:00 adaq Runsummary_Left_HRS
  • 70717 08/14/01 06:00 adaq Runsummary_Right_HRS
  • 70716 08/14/01 03:21 X. Zheng Target - EPR - 30.5% and New Conditions Aug-13
  • 70715 08/14/01 01:59 X. Zheng Laser alignment & power check
  • Monday

  • 70714 08/13/01 16:10 jpchen Target test: loss/temperature studies
  • 70713 08/13/01 13:06 jpchen Xiaofeng gets BeO viewer working
  • 70712 08/13/01 11:09 Bogdan/jpche Switched left gas Cherenkov #9 PMT. HVs set back for all channle for left Gas Cherenkov.
  • 70711 08/13/01 10:17 jpchen HVs for left gas Cherenkov turned off
  • 70710 08/13/01 10:15 jpchen spectrometers are set for next kinematics
  • 70708 08/13/01 07:37 kramerk Hall A in Restricted Access
  • 70707 08/13/01 07:23 kramerk Shift Summary
  • 70706 08/13/01 06:00 adaq Runsummary_Left_HRS
  • 70705 08/13/01 06:00 adaq Runsummary_Right_HRS
  • 70702 08/13/01 04:29 kramerk Beam delivery ended
  • 70700 08/13/01 03:53 apar Start_of_Run_Parity_1952
  • 70696 08/13/01 03:16 kramerk Several bad Start of Run headers
  • 70693 08/13/01 03:11 apar Start_of_Run_Parity_1951
  • 70690 08/13/01 03:05 kramerk Target - NMR/EPR - 44.7%/45.2%
  • 70689 08/13/01 02:41 kramerk Parity end of run failed
  • 70688 08/13/01 02:40 kramerk Beam check out
  • 70687 08/13/01 02:31 apar Start_of_Run_Parity_1950
  • 70682 08/13/01 02:08 apar Start_of_Run_Parity_1949
  • 70678 08/13/01 01:51 kramerk MCC sending pulsed beam!
  • Sunday

  • 70676 08/12/01 23:54 Armstrong shift summary :-(
  • 70675 08/12/01 20:09 Armstrong update from MCC
  • 70674 08/12/01 18:58 kramerk Target - NMR/EPR - 41.4%/42.3%
  • 70673 08/12/01 16:48 Armstrong Several magnet trips
  • 70671 08/12/01 16:01 adaq Day shift summary
  • 70668 08/12/01 12:24 W. Korsch CHL trip
  • 70667 08/12/01 11:54 W. Korsch Cosmics run
  • 70664 08/12/01 11:10 W. Korsch Target - EPR: 36.6%
  • 70663 08/12/01 11:08 T. Horn NMR
  • 70661 08/12/01 10:34 W. Korsch Accelerator lost RF
  • 70659 08/12/01 10:19 apar Start_of_Run_Parity_1948
  • 70655 08/12/01 09:38 apar Start_of_Run_Parity_1947
  • 70650 08/12/01 08:55 apar Start_of_Run_Parity_1946
  • 70644 08/12/01 08:13 apar Start_of_Run_Parity_1945
  • 70639 08/12/01 08:01 Zolnierczuk owl shift summary
  • 70634 08/12/01 07:30 X. Zheng Target - NMR 35.75%, EPR 37.16%
  • 70627 08/12/01 06:00 adaq Runsummary_Left_HRS
  • 70626 08/12/01 06:00 adaq Runsummary_Right_HRS
  • 70624 08/12/01 05:38 apar Start_of_Run_Parity_1942
  • 70618 08/12/01 04:57 apar Start_of_Run_Parity_1941
  • 70613 08/12/01 04:16 apar Start_of_Run_Parity_1940
  • 70609 08/12/01 03:24 apar Start_of_Run_Parity_1939
  • 70606 08/12/01 02:42 apar Start_of_Run_Parity_1938
  • 70601 08/12/01 02:34 apar Start_of_Run_Parity_1937
  • 70596 08/12/01 01:53 apar Start_of_Run_Parity_1936
  • 70592 08/12/01 01:19 Xiaofeng, Yu Target EPR: 37.8%
  • 70590 08/12/01 01:13 apar Start_of_Run_Parity_1935
  • 70588 08/12/01 00:57 Zolnierczuk target NMR P=36.5%
  • 70585 08/12/01 00:06 Armstrong swing shift run summary
  • Saturday

  • 70581 08/11/01 23:55 apar Start_of_Run_Parity_1934
  • 70576 08/11/01 23:09 apar Start_of_Run_Parity_1933
  • 70573 08/11/01 22:58 Armstrong beam current back to 12 muA
  • 70572 08/11/01 22:54 Armstrong beam current a bit low
  • 70568 08/11/01 22:28 apar Start_of_Run_Parity_1932
  • 70566 08/11/01 21:47 apar Start_of_Run_Parity_1931
  • 70560 08/11/01 21:05 apar Start_of_Run_Parity_1930
  • 70553 08/11/01 20:24 apar Start_of_Run_Parity_1929
  • 70551 08/11/01 20:08 Armstrong are there reference histograms?
  • 70550 08/11/01 19:45 Armstrong beam back
  • 70547 08/11/01 19:44 apar Start_of_Run_Parity_1928
  • 70542 08/11/01 19:24 apar Start_of_Run_Parity_1927
  • 70538 08/11/01 19:19 Armstrong Polarimetry
  • 70537 08/11/01 18:49 Xiaofeng, Co Target NMR:36.8% EPR:39.1%
  • 70533 08/11/01 17:28 Armstrong will take cosmics
  • 70530 08/11/01 17:25 apar Start_of_Run_Parity_1926
  • 70529 08/11/01 16:59 Armstrong go to longitudinal pumping
  • 70528 08/11/01 16:04 W. Korsch Day shift summary
  • 70524 08/11/01 14:43 W. Korsch Target - EPR: 36.3%
  • 70523 08/11/01 14:43 Sulkosky NMR polarization @ 35.0 %
  • 70520 08/11/01 14:35 apar Start_of_Run_Parity_1925
  • 70515 08/11/01 13:25 apar Start_of_Run_Parity_1924
  • 70512 08/11/01 13:02 Binet Compton HV
  • 70508 08/11/01 12:45 apar Start_of_Run_Parity_1923
  • 70506 08/11/01 12:26 W. Korsch Compton HV alarm
  • 70502 08/11/01 12:05 apar Start_of_Run_Parity_1922
  • 70497 08/11/01 11:24 apar Start_of_Run_Parity_1921
  • 70493 08/11/01 10:42 apar Start_of_Run_Parity_1920
  • 70488 08/11/01 10:02 apar Start_of_Run_Parity_1919
  • 70482 08/11/01 09:22 apar Start_of_Run_Parity_1918
  • 70480 08/11/01 08:55 Adaq NMR/EPR Polarization: 35.65; 36.8 %
  • 70476 08/11/01 08:41 apar Start_of_Run_Parity_1917
  • 70473 08/11/01 08:01 Zolnierczuk owl shift summary
  • 70469 08/11/01 07:32 apar Start_of_Run_Parity_1916
  • 70467 08/11/01 06:51 apar Start_of_Run_Parity_1915
  • 70461 08/11/01 06:10 apar Start_of_Run_Parity_1914
  • 70457 08/11/01 06:00 adaq Runsummary_Left_HRS
  • 70456 08/11/01 06:00 adaq Runsummary_Right_HRS
  • 70455 08/11/01 05:29 apar Start_of_Run_Parity_1913
  • 70449 08/11/01 04:45 apar Start_of_Run_Parity_1912
  • 70445 08/11/01 04:16 X. Zheng Target - NMR 36.11%, EPR 37.4%
  • 70443 08/11/01 04:05 apar Start_of_Run_Parity_1911
  • 70438 08/11/01 02:58 apar Start_of_Run_Parity_1910
  • 70433 08/11/01 02:17 apar Start_of_Run_Parity_1909
  • 70428 08/11/01 01:37 apar Start_of_Run_Parity_1908
  • 70422 08/11/01 00:56 apar Start_of_Run_Parity_1907
  • 70418 08/11/01 00:15 apar Start_of_Run_Parity_1906
  • 70413 08/11/01 00:01 finn shift summary
  • Friday

  • 70410 08/10/01 23:15 Xiaofeng, J. target NMR:36.8% EPR: 37.8%
  • 70406 08/10/01 22:39 apar Start_of_Run_Parity_1904
  • 70396 08/10/01 21:16 apar Start_of_Run_Parity_1902
  • 70387 08/10/01 19:49 apar Start_of_Run_Parity_1900
  • 70382 08/10/01 19:07 apar Start_of_Run_Parity_1899
  • 70377 08/10/01 18:26 Xiaofeng, J. Target NMR: 38% EPR: 38.4%
  • 70375 08/10/01 18:26 apar Start_of_Run_Parity_1898
  • 70371 08/10/01 17:33 apar Start_of_Run_Parity_1897
  • 70369 08/10/01 16:31 O. Hansen work4 cleanup
  • 70367 08/10/01 15:59 O. Hansen Shift Summary
  • 70366 08/10/01 15:31 averett Target at 90 degrees
  • 70365 08/10/01 15:28 averett NMR 37.2%
  • 70360 08/10/01 15:06 apar Start_of_Run_Parity_1896
  • 70355 08/10/01 14:24 apar Start_of_Run_Parity_1895
  • 70352 08/10/01 13:32 averett target at 270 again
  • 70351 08/10/01 13:29 W. Korsch Target - NMR: 37.3%, EPR: 38.7%
  • 70350 08/10/01 13:26 averett current ramping
  • 70347 08/10/01 13:22 apar Start_of_Run_Parity_1894
  • 70346 08/10/01 13:14 A.Vacheret charge asymmetry slope calibration
  • 70341 08/10/01 12:55 apar Start_of_Run_Parity_1893
  • 70337 08/10/01 12:47 apar Start_of_Run_Parity_1892
  • 70331 08/10/01 12:39 apar Start_of_Run_Parity_1891
  • 70327 08/10/01 12:32 apar Start_of_Run_Parity_1890
  • 70320 08/10/01 12:23 apar Start_of_Run_Parity_1889
  • 70315 08/10/01 12:18 apar Start_of_Run_Parity_1888
  • 70311 08/10/01 12:11 apar Start_of_Run_Parity_1887
  • 70306 08/10/01 11:33 apar Start_of_Run_Parity_1886
  • 70305 08/10/01 11:17 A. Deur eP trial.
  • 70300 08/10/01 10:53 apar Start_of_Run_Parity_1885
  • 70297 08/10/01 10:12 apar Start_of_Run_Parity_1884
  • 70289 08/10/01 09:08 apar Start_of_Run_Parity_1883
  • 70286 08/10/01 08:59 R. Michaels CODA problems last night
  • 70284 08/10/01 08:27 apar Start_of_Run_Parity_1882
  • 70280 08/10/01 08:15 Zolnierczuk owl shift summary
  • 70277 08/10/01 07:44 apar Start_of_Run_Parity_1881
  • 70274 08/10/01 07:03 apar Start_of_Run_Parity_1880
  • 70269 08/10/01 06:36 Zolnierczuk target EPR=36.7%
  • 70266 08/10/01 06:20 apar Start_of_Run_Parity_1879
  • 70263 08/10/01 06:00 adaq Runsummary_Left_HRS
  • 70262 08/10/01 06:00 adaq Runsummary_Right_HRS
  • 70260 08/10/01 05:39 apar Start_of_Run_Parity_1878
  • 70251 08/10/01 05:29 steiner Target NMR: 35.2%
  • 70238 08/10/01 03:38 apar Start_of_Run_Parity_1877
  • 70230 08/10/01 02:48 apar Start_of_Run_Parity_1876
  • 70226 08/10/01 02:07 apar Start_of_Run_Parity_1875
  • 70222 08/10/01 01:26 apar Start_of_Run_Parity_1874
  • 70217 08/10/01 00:45 apar Start_of_Run_Parity_1873
  • 70212 08/10/01 00:05 xfzhu, J. Ro Target NMR:35.8% EPR:38.3%
  • 70209 08/10/01 00:03 apar Start_of_Run_Parity_1872
  • 70207 08/10/01 00:02 Finn shift summary
  • Thursday

  • 70202 08/09/01 22:47 apar Start_of_Run_Parity_1871
  • 70197 08/09/01 22:06 apar Start_of_Run_Parity_1870
  • 70192 08/09/01 21:26 apar Start_of_Run_Parity_1869
  • 70189 08/09/01 20:47 finn Beam half wave plate OUT
  • 70188 08/09/01 20:45 apar Start_of_Run_Parity_1868
  • 70183 08/09/01 20:21 finn beam half wave plate put IN for next half hour
  • 70182 08/09/01 20:15 apar Start_of_Run_Parity_1867
  • 70175 08/09/01 19:36 apar Start_of_Run_Parity_1866
  • 70171 08/09/01 19:05 apar Start_of_Run_Parity_1865
  • 70167 08/09/01 18:51 xfzhu problem with cha2hp printer
  • 70166 08/09/01 18:39 averett beam position
  • 70165 08/09/01 18:33 xfzhu, J. Ro target NMR:36.8% EPR:38.3%
  • 70162 08/09/01 18:24 apar Start_of_Run_Parity_1864
  • 70157 08/09/01 17:24 apar Start_of_Run_Parity_1863
  • 70153 08/09/01 16:43 apar Start_of_Run_Parity_1862
  • 70149 08/09/01 16:41 xfzhu scaler update
  • 70148 08/09/01 16:32 milbrath beam spot off
  • 70145 08/09/01 16:02 apar Start_of_Run_Parity_1861
  • 70141 08/09/01 15:53 averett Day Shift Summary
  • 70140 08/09/01 15:23 Saha Arc Integral Measurement
  • 70137 08/09/01 15:18 apar Start_of_Run_Parity_1860
  • 70134 08/09/01 15:08 W. Korsch Target Spin at 270 deg
  • 70133 08/09/01 14:45 X. Zheng, W. Target - NMR: 38.2%, EPR:
  • 70130 08/09/01 14:37 apar Start_of_Run_Parity_1859
  • 70127 08/09/01 14:00 xfzhu scaler update
  • 70125 08/09/01 13:57 apar Start_of_Run_Parity_1858
  • 70119 08/09/01 13:16 apar Start_of_Run_Parity_1857
  • 70118 08/09/01 13:13 A. Deur eP run (not very good)
  • 70117 08/09/01 12:51 averett beam half wave plate OUT
  • 70112 08/09/01 11:40 apar Start_of_Run_Parity_1856
  • 70109 08/09/01 11:20 apar Start_of_Run_Parity_1855
  • 70104 08/09/01 10:55 apar Start_of_Run_Parity_1854
  • 70101 08/09/01 09:31 W. Korsch Target - NMR: 35.3%, EPR: 37.0%
  • 70098 08/09/01 08:30 averett MEDM terminal dying
  • 70094 08/09/01 08:07 apar Start_of_Run_Parity_1853
  • 70092 08/09/01 08:06 kramerk Shift Summary
  • 70088 08/09/01 07:26 apar Start_of_Run_Parity_1852
  • 70083 08/09/01 06:45 apar Start_of_Run_Parity_1851
  • 70078 08/09/01 06:04 apar Start_of_Run_Parity_1850
  • 70075 08/09/01 06:00 adaq Runsummary_Left_HRS
  • 70074 08/09/01 06:00 adaq Runsummary_Right_HRS
  • 70071 08/09/01 05:23 apar Start_of_Run_Parity_1849
  • 70070 08/09/01 05:22 kramerk MCC says they are ready to send beam
  • 70067 08/09/01 05:10 kramerk Another RF Zone down
  • 70064 08/09/01 04:57 apar Start_of_Run_Parity_1848
  • 70058 08/09/01 04:17 apar Start_of_Run_Parity_1847
  • 70056 08/09/01 03:53 kramerk MCC says they have recovered
  • 70055 08/09/01 03:51 kramerk MCC monitor in halla counting house is purple
  • 70054 08/09/01 03:49 kramerk Target - NMR/EPR - 35.3%/37.1%
  • 70053 08/09/01 03:45 kramerk MCC says beam down for a few minutes
  • 70050 08/09/01 03:37 apar Start_of_Run_Parity_1846
  • 70044 08/09/01 02:32 apar Start_of_Run_Parity_1845
  • 70039 08/09/01 01:52 apar Start_of_Run_Parity_1844
  • 70038 08/09/01 01:51 kramerk MCC says beam is coming back!
  • 70037 08/09/01 01:37 kramerk MCC requests time for QE measurement
  • 70033 08/09/01 01:23 apar Start_of_Run_Parity_1843
  • 70029 08/09/01 00:42 apar Start_of_Run_Parity_1842
  • 70025 08/09/01 00:05 jpchen swing shift summary
  • 70020 08/09/01 00:01 apar Start_of_Run_Parity_1841
  • Wednesday

  • 70016 08/08/01 23:21 apar Start_of_Run_Parity_1840
  • 70012 08/08/01 23:12 apar Start_of_Run_Parity_1839
  • 70005 08/08/01 22:09 apar Start_of_Run_Parity_1838
  • 70003 08/08/01 21:52 jpchen Longitudinal and transverse temperature difference
  • 70002 08/08/01 21:41 jpchen Target: EPR=37.8%
  • 69998 08/08/01 21:28 apar Start_of_Run_Parity_1837
  • 69997 08/08/01 21:08 Olivier NMR measurement
  • 69993 08/08/01 20:13 apar Start_of_Run_Parity_1836
  • 69988 08/08/01 19:33 apar Start_of_Run_Parity_1835
  • 69983 08/08/01 18:52 apar Start_of_Run_Parity_1834
  • 69978 08/08/01 18:11 apar Start_of_Run_Parity_1833
  • 69973 08/08/01 17:29 apar Start_of_Run_Parity_1832
  • 69968 08/08/01 16:48 apar Start_of_Run_Parity_1831
  • 69962 08/08/01 16:08 R. Michaels day shift summary
  • 69960 08/08/01 16:07 apar Start_of_Run_Parity_1830
  • 69958 08/08/01 15:46 W. Korsch Target - NMR: 35.6%, EPR: 37.8%
  • 69957 08/08/01 15:41 Zolnierczuk runstart/end scripts fixed
  • 69953 08/08/01 15:26 apar Start_of_Run_Parity_1829
  • 69952 08/08/01 15:20 R. Michaels slightly corrected figure (chg asy)
  • 69950 08/08/01 14:20 R.Michaels roc15 hiccup
  • 69948 08/08/01 14:19 apar Start_of_Run_Parity_1828
  • 69946 08/08/01 14:15 R. Michaels charge asmmetry in 3 days
  • 69942 08/08/01 13:33 apar Start_of_Run_Parity_1827
  • 69936 08/08/01 12:52 apar Start_of_Run_Parity_1826
  • 69934 08/08/01 12:27 R. Michaels hbook's written to work4 now
  • 69933 08/08/01 12:12 apar Start_of_Run_Parity_1825
  • 69927 08/08/01 12:04 W. Korsch Target - NMR: 36.4%, EPR: 37.8 %
  • 69926 08/08/01 11:52 R. Michaels target halfwave IN
  • 69925 08/08/01 11:47 W. Korsch B-field at 90 degrees
  • 69922 08/08/01 11:24 apar Start_of_Run_Parity_1824
  • 69921 08/08/01 11:11 K. McCormick EPICS changes started
  • 69917 08/08/01 10:31 apar Start_of_Run_Parity_1823
  • 69912 08/08/01 09:41 apar Start_of_Run_Parity_1822
  • 69906 08/08/01 08:51 apar Start_of_Run_Parity_1821
  • 69904 08/08/01 08:04 kramerk Shift Summary
  • 69900 08/08/01 07:48 apar Start_of_Run_Parity_1820
  • 69895 08/08/01 06:58 apar Start_of_Run_Parity_1819
  • 69890 08/08/01 06:09 apar Start_of_Run_Parity_1818
  • 69889 08/08/01 06:00 adaq Runsummary_Left_HRS
  • 69888 08/08/01 06:00 adaq Runsummary_Right_HRS
  • 69887 08/08/01 05:48 kramerk Target - NMR/EPR - 37.2%/39.0%
  • 69886 08/08/01 05:37 kramerk Target moved to Pol3he, transverse lasers on
  • 69885 08/08/01 05:33 kramerk Raster Check
  • 69880 08/08/01 05:29 apar Start_of_Run_Parity_1817
  • 69879 08/08/01 03:52 kramerk Target - Field rotated to 0 for temporary pumping
  • 69876 08/08/01 03:30 apar Start_of_Run_Parity_1816
  • 69869 08/08/01 02:40 apar Start_of_Run_Parity_1815
  • 69864 08/08/01 01:50 apar Start_of_Run_Parity_1814
  • 69861 08/08/01 01:13 kramerk Target - NMR/EPR - 36.7%/38.5%
  • 69858 08/08/01 01:01 apar Start_of_Run_Parity_1813
  • 69857 08/08/01 01:00 apar Start_of_Run_Parity_1812
  • 69853 08/08/01 00:04 J. Roche shift summary
  • Tuesday

  • 69851 08/07/01 23:41 apar Start_of_Run_Parity_1811
  • 69847 08/07/01 22:54 Olivier adaql2 crash, parity runs screwed
  • 69844 08/07/01 22:51 apar Start_of_Run_Parity_1810
  • 69843 08/07/01 22:50 apar Start_of_Run_Parity_1809
  • 69842 08/07/01 22:49 apar Start_of_Run_Parity_1808
  • 69841 08/07/01 22:48 apar Start_of_Run_Parity_1807
  • 69836 08/07/01 22:46 apar Start_of_Run_Parity_1806
  • 69831 08/07/01 22:42 apar Start_of_Run_Parity_1805
  • 69822 08/07/01 22:02 apar Start_of_Run_Parity_1806
  • 69818 08/07/01 21:52 A. Deur Elastic asymmetry for both arm
  • 69815 08/07/01 21:07 apar Start_of_Run_Parity_1805
  • 69813 08/07/01 20:27 kramerk Target - NMR/EPR - 37.6%/39.3%
  • 69812 08/07/01 20:25 busato junk runs : 2671/21671/2672/21672/2673/21673
  • 69809 08/07/01 20:17 apar Start_of_Run_Parity_1804
  • 69803 08/07/01 20:12 apar Start_of_Run_Parity_1803
  • 69800 08/07/01 20:02 apar Start_of_Run_Parity_1802
  • 69797 08/07/01 19:58 apar Start_of_Run_Parity_1801
  • 69792 08/07/01 19:04 apar Start_of_Run_Parity_1800
  • 69787 08/07/01 18:15 apar Start_of_Run_Parity_1799
  • 69786 08/07/01 18:14 averett Empty target position changed
  • 69785 08/07/01 17:56 Wolfgang/jpc back to transvers pumping *(late entry*)
  • 69784 08/07/01 17:52 wolfgang/jpc Target, NMR=39.1%, EPR=40.8%
  • 69783 08/07/01 17:41 averett new target encoder settings
  • 69780 08/07/01 17:36 busato lot of trips for run 2667.
  • 69779 08/07/01 17:36 averett new beam/target position
  • 69776 08/07/01 17:27 apar Start_of_Run_Parity_1798
  • 69773 08/07/01 17:13 Binet Slow Controls of the Compton Polarimeter
  • 69770 08/07/01 16:36 apar Start_of_Run_Parity_1797
  • 69765 08/07/01 16:27 apar Start_of_Run_Parity_1796
  • 69760 08/07/01 16:19 apar Start_of_Run_Parity_1795
  • 69759 08/07/01 16:00 R. Michaels day shift summary
  • 69754 08/07/01 15:19 averett left-arm scintillator 6 left side
  • 69753 08/07/01 15:11 Antonin little modification in charge asym fast feedback software.
  • 69750 08/07/01 15:09 apar Start_of_Run_Parity_1794
  • 69746 08/07/01 14:19 apar Start_of_Run_Parity_1793
  • 69741 08/07/01 13:29 apar Start_of_Run_Parity_1792
  • 69739 08/07/01 11:52 T. Averett, Target -NMR: 40.6%, EPR: 41.5%
  • 69738 08/07/01 11:48 averett R-arm cerenkov HV restored
  • 69737 08/07/01 11:45 A. Deur Elastic asymmetry, left arm
  • 69736 08/07/01 11:10 averett right cherenkov high voltages
  • 69732 08/07/01 10:54 apar Start_of_Run_Parity_1791
  • 69728 08/07/01 10:40 xfzhu junk files in espace/hbook directory
  • 69727 08/07/01 10:27 averett cerenkov high voltages
  • 69726 08/07/01 10:11 R. Michaels gas cherenkov gains low (R-arm)
  • 69724 08/07/01 10:01 apar Start_of_Run_Parity_1790
  • 69721 08/07/01 09:54 averett current not ramping after trip
  • 69720 08/07/01 09:37 R. Michaels preshower block 64
  • 69717 08/07/01 09:18 R. Michaels MCC had chg asy lock off last night
  • 69712 08/07/01 09:10 apar Start_of_Run_Parity_1789
  • 69711 08/07/01 09:02 xfzhu false asymmetry from scaler
  • 69708 08/07/01 08:19 apar Start_of_Run_Parity_1788
  • 69707 08/07/01 08:10 averett OTR no raster
  • 69706 08/07/01 08:08 averett Harp scan
  • 69704 08/07/01 08:00 milbrath owl shift summary
  • 69698 08/07/01 07:38 milbrath holding field change
  • 69688 08/07/01 06:33 apar Start_of_Run_Parity_1785
  • 69685 08/07/01 06:28 apar Start_of_Run_Parity_1784
  • 0 08/07/01 06:28 no_name no_key
  • 69682 08/07/01 06:00 adaq Runsummary_Left_HRS
  • 69681 08/07/01 06:00 adaq Runsummary_Right_HRS
  • 69680 08/07/01 05:56 kramerk Target - NMR/EPR - 44.9%/46.0%
  • 69679 08/07/01 05:30 milbrath no beam
  • 69678 08/07/01 05:28 milbrath UPS
  • 69674 08/07/01 02:36 apar Start_of_Run_Parity_1783
  • 69670 08/07/01 02:09 kramerk Target - NMR/EPR - 43.9%/46.2%
  • 69669 08/07/01 01:59 kramerk DAQ - Rebooted adaqcp and ROC14
  • 69667 08/07/01 01:44 apar Start_of_Run_Parity_1782
  • 69662 08/07/01 00:35 xfzhu scalers' result till run2644
  • 69659 08/07/01 00:13 apar Start_of_Run_Parity_1781
  • 69658 08/07/01 00:13 O. Hansen adaql2 kernel upgrade
  • 69654 08/07/01 00:03 O. Hansen Swing shift summary
  • Monday

  • 69653 08/06/01 23:57 O. Hansen Run numbers out of sync
  • 69652 08/06/01 23:25 xfzhu root file for scalers
  • 69651 08/06/01 23:24 apar Start_of_Run_Parity_1780
  • 69644 08/06/01 22:35 apar Start_of_Run_Parity_1779
  • 69641 08/06/01 21:46 apar Start_of_Run_Parity_1778
  • 69640 08/06/01 21:46 A. Deur He3 pressure curve
  • 69639 08/06/01 21:38 O. Hansen Half-wave plate IN
  • 69638 08/06/01 21:33 W. Korsch Target - EPR - 46.9%
  • 69637 08/06/01 21:31 Olivier NMR measurement
  • 69633 08/06/01 21:10 Eugene Moller results
  • 69630 08/06/01 21:04 apar Start_of_Run_Parity_1777
  • 69628 08/06/01 20:54 olivier harp scan
  • 69627 08/06/01 18:34 A. Deur printcharge
  • 69626 08/06/01 17:40 W. Korsch Target - EPR: 47.1%
  • 69625 08/06/01 17:19 R. Michaels charge asymmetry feedback results
  • 69622 08/06/01 16:22 apar Start_of_Run_Parity_1776
  • 69618 08/06/01 16:13 jpchen day shift summary
  • 69617 08/06/01 16:12 averett OTR at 3.466 GeV
  • 69615 08/06/01 15:45 jpchen Current raised to 12 uA
  • 69614 08/06/01 15:42 R. Michaels wrong halog EndRun info - fixed
  • 69613 08/06/01 15:35 W. Korsch What to do when "Crate Resets" screen fails!
  • 69611 08/06/01 15:32 apar Start_of_Run_Parity_1775
  • 69605 08/06/01 14:27 apar Start_of_Run_Parity_1774
  • 69601 08/06/01 13:52 Reitz 2641/21641 is junk
  • 69600 08/06/01 13:51 Reitz DAQ,EndOfRun 2640, ROC14
  • 69598 08/06/01 13:37 apar Start_of_Run_Parity_1773
  • 69594 08/06/01 12:53 R. Michaels Recovery from adaql2 crash
  • 69593 08/06/01 12:47 jpchen target: NMR=46.7%%, EPR=48.7%
  • 69590 08/06/01 12:41 apar Start_of_Run_Parity_1772
  • 69586 08/06/01 12:37 apar Start_of_Run_Parity_1771
  • 69584 08/06/01 12:33 apar Start_of_Run_Parity_1765
  • 69581 08/06/01 12:15 R. Michaels adaql2 frozen, reboot
  • 69579 08/06/01 11:55 W. Korsch Shower/Preshower ADC spectra
  • 69578 08/06/01 11:52 A. Deur, W. Pions and Electrons for L-HRS
  • 69576 08/06/01 11:41 apar Start_of_Run_Parity_1768
  • 69570 08/06/01 10:50 apar Start_of_Run_Parity_1767
  • 69566 08/06/01 10:30 jpchen Left arn ADC S1_P2 extra hits
  • 69565 08/06/01 10:20 R. Michaels Extra ADC hit "problem"
  • 69564 08/06/01 10:04 jpchen Reference cell with N2 also leaks
  • 69559 08/06/01 09:32 R. Michaels event sizes ok
  • 69556 08/06/01 09:08 apar Start_of_Run_Parity_1765
  • 69552 08/06/01 08:19 apar Start_of_Run_Parity_1764
  • 69545 08/06/01 08:10 kramerk Shift Summary
  • 69542 08/06/01 07:58 apar Start_of_Run_Parity_1763
  • 69535 08/06/01 07:53 apar Start_of_Run_Parity_1762
  • 69534 08/06/01 07:45 kramerk Beam spot finally acceptable
  • 69533 08/06/01 07:36 kramerk Target - NMR/EPR - 49.2%/50.3%
  • 69532 08/06/01 06:55 kramerk MCC still trying to fix beam size problem
  • 69531 08/06/01 06:00 adaq Runsummary_Left_HRS
  • 69530 08/06/01 06:00 adaq Runsummary_Right_HRS
  • 69527 08/06/01 04:13 apar Start_of_Run_Parity_1761
  • 69521 08/06/01 04:03 kramerk Beam size still too small, MCC bringing in expert
  • 69518 08/06/01 03:17 apar Start_of_Run_Parity_1760
  • 69516 08/06/01 03:13 A. Deur N2 contamination of the He3 ref cell.
  • 69515 08/06/01 03:10 kramerk Moller quads not at polarization values
  • 69514 08/06/01 02:50 kramerk Harp Scan 1362
  • 69513 08/06/01 02:48 apar Start_of_Run_Parity_1759
  • 69509 08/06/01 02:36 kramerk Harp scan 1361
  • 69508 08/06/01 00:02 J. Roche evening Shift summary
  • Sunday

  • 69506 08/05/01 23:39 A. Deur nice N2 pressure curves
  • 69502 08/05/01 22:54 ransome event size left and right different
  • 69501 08/05/01 22:16 xfzhu Target, NMR:51.62% EPR: 48.9%
  • 69500 08/05/01 20:39 J. Roche beam update
  • 69498 08/05/01 20:30 apar Start_of_Run_Parity_1758
  • 69496 08/05/01 20:19 xfzhu Left arm S1 ADC hit pattern
  • 69493 08/05/01 19:45 apar Start_of_Run_Parity_1757
  • 69485 08/05/01 18:58 apar Start_of_Run_Parity_1756
  • 69478 08/05/01 18:36 apar Start_of_Run_Parity_1755
  • 69473 08/05/01 18:25 apar Start_of_Run_Parity_1754
  • 69469 08/05/01 17:07 kramerk Target - NMR @ 15:45 - 50.8%
  • 69466 08/05/01 16:49 apar Start_of_Run_Parity_1753
  • 69462 08/05/01 16:32 apar Start_of_Run_Parity_1752
  • 69457 08/05/01 16:14 apar Start_of_Run_Parity_1751
  • 69452 08/05/01 16:02 ransome day shift summary
  • 69448 08/05/01 15:48 apar Start_of_Run_Parity_1750
  • 69443 08/05/01 14:22 apar Start_of_Run_Parity_1749
  • 69441 08/05/01 13:58 ransome Cerenkov hit distribution
  • 69437 08/05/01 13:46 apar Start_of_Run_Parity_1748
  • 69436 08/05/01 13:30 ransome tkABU problem
  • 69435 08/05/01 13:27 ransome dead time problem
  • 69433 08/05/01 13:24 X. Zheng Target - EPR conditions
  • 69430 08/05/01 13:18 apar Start_of_Run_Parity_1747
  • 69424 08/05/01 12:55 apar Start_of_Run_Parity_1746
  • 69421 08/05/01 12:53 W. Korsch Late entry: Network/DAQ problems
  • 69419 08/05/01 12:36 apar Start_of_Run_Parity_1745
  • 69415 08/05/01 12:21 W. Korsch Target - EPR: 49%
  • 69414 08/05/01 12:17 Adaq S1/S2 ADC/TDC changed from Reference
  • 69412 08/05/01 12:15 apar Start_of_Run_Parity_1744
  • 69406 08/05/01 11:49 apar Start_of_Run_Parity_1743
  • 69403 08/05/01 11:39 R. Michaels R-arm CODA problems, solution
  • 69402 08/05/01 11:36 apar Start_of_Run_Parity_1742
  • 69393 08/05/01 11:08 apar Start_of_Run_Parity_1741
  • 69388 08/05/01 11:01 apar Start_of_Run_Parity_1740
  • 69384 08/05/01 10:35 apar Start_of_Run_Parity_1739
  • 69378 08/05/01 10:24 apar Start_of_Run_Parity_1738
  • 69375 08/05/01 09:50 apar Start_of_Run_Parity_1737
  • 69372 08/05/01 09:43 A. Deur eP energy measurement: 1.198 GeV
  • 69368 08/05/01 08:04 Sulkosky Owl shift Summary
  • 69367 08/05/01 08:00 apar Start_of_Run_Parity_1736
  • 69363 08/05/01 07:20 sulkosky Computer problems.
  • 69357 08/05/01 06:35 apar Start_of_Run_Parity_1734
  • 69352 08/05/01 06:24 apar Start_of_Run_Parity_1733
  • 69348 08/05/01 06:20 apar Start_of_Run_Parity_1732
  • 69343 08/05/01 06:15 apar Start_of_Run_Parity_1731
  • 69338 08/05/01 06:10 apar Start_of_Run_Parity_1730
  • 69334 08/05/01 06:00 adaq Runsummary_Left_HRS
  • 69333 08/05/01 06:00 adaq Runsummary_Right_HRS
  • 69331 08/05/01 05:53 apar Start_of_Run_Parity_1729
  • 69327 08/05/01 05:49 sulkosky Momentum change.
  • 69326 08/05/01 05:49 apar Start_of_Run_Parity_1728
  • 69320 08/05/01 05:08 xfzhu Target EPR: NMR:48.03%
  • 69317 08/05/01 04:41 Sulkosky Beam position.
  • 69316 08/05/01 04:41 apar Start_of_Run_Parity_1727
  • 69314 08/05/01 04:29 sulkosky CODA problems
  • 69313 08/05/01 04:19 Sulkosky MCC cycling magnets.
  • 69309 08/05/01 03:57 Sulkosky Beam Back!
  • 69301 08/05/01 02:45 Sulkosky MCC lost a Magnet.
  • 69293 08/05/01 01:58 apar Start_of_Run_Parity_1723
  • 69289 08/05/01 01:30 apar Start_of_Run_Parity_1722
  • 69282 08/05/01 01:13 apar Start_of_Run_Parity_1721
  • 69277 08/05/01 00:46 apar Start_of_Run_Parity_1720
  • 69270 08/05/01 00:16 apar Start_of_Run_Parity_1719
  • 69269 08/05/01 00:06 xfzhu Asymmetry from scalers
  • Saturday

  • 69268 08/04/01 23:58 A. Deur Beautiful N2 peaks
  • 69267 08/04/01 23:58 Finn Shift summary
  • 69265 08/04/01 23:48 apar Start_of_Run_Parity_1718
  • 69259 08/04/01 23:19 apar Start_of_Run_Parity_1717
  • 69255 08/04/01 22:49 J. Roche, W. Target - NMR: 49.4%, EPR: 48.0%
  • 69254 08/04/01 22:45 finn epr polarization is 48%
  • 69253 08/04/01 22:40 finn complete target pol measurements, change beam lambda plate to IN position
  • 69251 08/04/01 22:39 apar Start_of_Run_Parity_1716
  • 69249 08/04/01 22:19 finn beam off for ep measurement
  • 69244 08/04/01 21:48 apar Start_of_Run_Parity_1715
  • 69241 08/04/01 21:06 apar Start_of_Run_Parity_1714
  • 69236 08/04/01 20:36 apar Start_of_Run_Parity_1713
  • 69230 08/04/01 19:50 apar Start_of_Run_Parity_1712
  • 69225 08/04/01 19:24 finn run 2574 with 6 mstr collimator
  • 69224 08/04/01 19:16 A. Deur test run with collimator in
  • 69223 08/04/01 19:15 apar Start_of_Run_Parity_1711
  • 69218 08/04/01 18:44 apar Start_of_Run_Parity_1710
  • 69216 08/04/01 18:43 apar Start_of_Run_Parity_1710
  • 69215 08/04/01 18:42 finn setting prescale factors to 2/2/2/2
  • 69211 08/04/01 18:32 apar Start_of_Run_Parity_1709
  • 69207 08/04/01 18:28 finn first spectra from shapiro
  • 69205 08/04/01 18:26 apar Start_of_Run_Parity_1708
  • 69200 08/04/01 17:59 apar Start_of_Run_Parity_1707
  • 69196 08/04/01 17:45 W. Korsch 3He bottle closed
  • 69193 08/04/01 17:43 apar Start_of_Run_Parity_1706
  • 69192 08/04/01 17:38 J. Roche rought check for he3 pressure curve
  • 69191 08/04/01 17:00 W. Korsch Reference Cell leaks for 3He
  • 69190 08/04/01 16:31 W. Korsch Target - EPR: 48.1%
  • 69185 08/04/01 16:07 apar Start_of_Run_Parity_1705
  • 69181 08/04/01 16:05 Ransome Shift summary
  • 69180 08/04/01 16:03 kramerk Target - NMR Calibration Constant
  • 69176 08/04/01 15:47 apar Start_of_Run_Parity_1704
  • 69174 08/04/01 15:38 kramerk Target - EPR taken @ 11:45 - 48.2%
  • 69172 08/04/01 15:27 Reitz Left Arm Helicity gates II
  • 69171 08/04/01 15:27 apar Start_of_Run_Parity_1703
  • 69165 08/04/01 15:14 Reitz Left Arm Helicity gates
  • 69159 08/04/01 14:19 apar Start_of_Run_Parity_1702
  • 69155 08/04/01 13:50 W. Korsch, B More on scaler offsets
  • 69153 08/04/01 13:48 apar Start_of_Run_Parity_1701
  • 69148 08/04/01 13:10 apar Start_of_Run_Parity_1700
  • 69144 08/04/01 12:56 Reitz Scalers and their sums
  • 69142 08/04/01 12:48 apar Start_of_Run_Parity_1699
  • 69137 08/04/01 12:26 apar Start_of_Run_Parity_1698
  • 69131 08/04/01 12:21 apar Start_of_Run_Parity_1697
  • 69127 08/04/01 11:50 apar Start_of_Run_Parity_1696
  • 69125 08/04/01 11:45 X. Zheng sample histogram not available yet
  • 69124 08/04/01 11:32 Ransome reference cell spectrum
  • 69119 08/04/01 10:59 apar Start_of_Run_Parity_1695
  • 69114 08/04/01 10:37 apar Start_of_Run_Parity_1694
  • 69113 08/04/01 10:27 xfzhu Target NMR(taken) EPR 48:6%
  • 69112 08/04/01 10:27 Tanja Horn raster for reference cell
  • 69107 08/04/01 10:06 apar Start_of_Run_Parity_1693
  • 69102 08/04/01 10:04 apar Start_of_Run_Parity_1692
  • 69097 08/04/01 09:41 apar Start_of_Run_Parity_1691
  • 69093 08/04/01 09:23 apar Start_of_Run_Parity_1690
  • 69090 08/04/01 09:14 kramerk Target - PC downstairs needs to be rebooted
  • 69088 08/04/01 09:13 apar Start_of_Run_Parity_1689
  • 69082 08/04/01 08:54 averett xscaler ok
  • 69080 08/04/01 08:47 apar Start_of_Run_Parity_1688
  • 69076 08/04/01 08:45 xfzhu Left arm scaler
  • 69075 08/04/01 08:37 W. Korsch Problems with BCM values
  • 69074 08/04/01 08:31 averett xscaler rates
  • 69072 08/04/01 08:30 apar Start_of_Run_Parity_1687
  • 69067 08/04/01 08:08 apar Start_of_Run_Parity_1686
  • 69062 08/04/01 08:01 xfzhu Fill reference cell with N2 and continue pumping
  • 69060 08/04/01 07:49 apar Start_of_Run_Parity_1685
  • 69053 08/04/01 07:32 apar Start_of_Run_Parity_1684
  • 69049 08/04/01 06:53 apar Start_of_Run_Parity_1683
  • 69044 08/04/01 06:35 apar Start_of_Run_Parity_1682
  • 69040 08/04/01 06:32 xfzhu scaler analysis for false aymmetry run( bema half wave plate in)
  • 69037 08/04/01 06:18 apar Start_of_Run_Parity_1681
  • 69033 08/04/01 06:02 apar Start_of_Run_Parity_1680
  • 69030 08/04/01 06:00 adaq Runsummary_Left_HRS
  • 69029 08/04/01 06:00 adaq Runsummary_Right_HRS
  • 69026 08/04/01 05:45 apar Start_of_Run_Parity_1679
  • 69021 08/04/01 05:26 apar Start_of_Run_Parity_1678
  • 69018 08/04/01 05:10 apar Start_of_Run_Parity_1677
  • 69015 08/04/01 05:09 apar Start_of_Run_Parity_1676
  • 69014 08/04/01 05:01 Guy Ron Half-Wave Plate
  • 69009 08/04/01 04:41 apar Start_of_Run_Parity_1675
  • 69003 08/04/01 04:23 apar Start_of_Run_Parity_1674
  • 68998 08/04/01 04:06 apar Start_of_Run_Parity_1673
  • 68994 08/04/01 03:47 apar Start_of_Run_Parity_1672
  • 68988 08/04/01 03:30 apar Start_of_Run_Parity_1671
  • 68985 08/04/01 03:25 xfzhu usage of "printcharge"
  • 68982 08/04/01 03:13 apar Start_of_Run_Parity_1670
  • 68980 08/04/01 03:03 B Milbrath charge asymmetry
  • 68977 08/04/01 02:56 apar Start_of_Run_Parity_1669
  • 68972 08/04/01 02:38 apar Start_of_Run_Parity_1668
  • 68969 08/04/01 02:36 R. Michaels feedback not working
  • 68966 08/04/01 02:28 apar Start_of_Run_Parity_1667
  • 68963 08/04/01 02:27 Xfzhu Reference cell N2 filled and pumping
  • 68954 08/04/01 02:09 apar Start_of_Run_Parity_1665
  • 68949 08/04/01 02:02 apar Start_of_Run_Parity_1664
  • 68946 08/04/01 01:51 apar Start_of_Run_Parity_1663
  • 68940 08/04/01 01:43 apar Start_of_Run_Parity_1662
  • 68937 08/04/01 01:33 apar Start_of_Run_Parity_1661
  • 68936 08/04/01 01:32 X. Zheng, W. Target - EPR - 44.9%
  • 68934 08/04/01 01:25 slifer Access
  • 68933 08/04/01 01:21 X. Zheng charges in end_of_run files
  • 68932 08/04/01 01:20 G Ron DAQ
  • 68928 08/04/01 01:01 apar Start_of_Run_Parity_1660
  • 68924 08/04/01 00:48 apar Start_of_Run_Parity_1659
  • 68923 08/04/01 00:31 G. Ron NMR polarization
  • Friday

  • 68921 08/03/01 23:49 slifer slifer
  • 68919 08/03/01 23:31 apar Start_of_Run_Parity_1658
  • 68914 08/03/01 23:14 slifer Right arm scintillators
  • 68913 08/03/01 23:10 finn datamon not working on left spectrometer
  • 68912 08/03/01 23:07 apar Start_of_Run_Parity_1657
  • 68910 08/03/01 23:07 X. Zheng EPR calibration constant for Shapiro with 3 lasers
  • 68907 08/03/01 23:02 slifer daq
  • 68906 08/03/01 23:02 apar Start_of_Run_Parity_1656
  • 68903 08/03/01 22:47 finn reboot ts1 after run 2514
  • 68900 08/03/01 22:46 apar Start_of_Run_Parity_1655
  • 68898 08/03/01 22:38 finn roc2 has multiple interupts
  • 68897 08/03/01 22:33 slifer Left arm End of Run file prob
  • 68894 08/03/01 22:25 apar Start_of_Run_Parity_1654
  • 68890 08/03/01 22:11 apar Start_of_Run_Parity_1653
  • 68884 08/03/01 21:55 apar Start_of_Run_Parity_1652
  • 68879 08/03/01 21:44 apar Start_of_Run_Parity_1651
  • 68875 08/03/01 21:33 finn DAQ problem run 2509
  • 68873 08/03/01 21:32 apar Start_of_Run_Parity_1650
  • 68869 08/03/01 21:23 apar Start_of_Run_Parity_1649
  • 68865 08/03/01 21:08 apar Start_of_Run_Parity_1648
  • 68861 08/03/01 20:59 finn right arm bottom HV
  • 68860 08/03/01 20:58 slifer daq
  • 68859 08/03/01 20:58 finn more scintillator high voltages
  • 68858 08/03/01 20:56 finn right arm top SCINTILLATORS
  • 68855 08/03/01 20:51 apar Start_of_Run_Parity_1647
  • 68852 08/03/01 20:50 Finn resynch run numbers
  • 68850 08/03/01 20:37 apar Start_of_Run_Parity_1646
  • 68848 08/03/01 20:35 apar Start_of_Run_Parity_1645
  • 68845 08/03/01 20:34 apar Start_of_Run_Parity_1644
  • 68841 08/03/01 19:51 apar Start_of_Run_Parity_1643
  • 68839 08/03/01 19:32 kramerk Target - Results of temperature study
  • 68838 08/03/01 19:18 Eugene Moller results
  • 68837 08/03/01 18:58 jpchen Reference cell pump/purge
  • 68836 08/03/01 17:23 slifer Right arm shower
  • 68833 08/03/01 16:57 slifer Right arm shower
  • 68831 08/03/01 16:56 apar Start_of_Run_Parity_1642
  • 68827 08/03/01 16:45 slifer Scintillator TDC's
  • 68825 08/03/01 16:43 apar Start_of_Run_Parity_1641
  • 68821 08/03/01 16:40 slifer beam offset
  • 68820 08/03/01 16:28 apar Start_of_Run_Parity_1640
  • 68815 08/03/01 16:05 apar Start_of_Run_Parity_1639
  • 68810 08/03/01 15:45 averett Day shift summary
  • 68807 08/03/01 15:44 apar Start_of_Run_Parity_1638
  • 68804 08/03/01 14:50 apar Start_of_Run_Parity_1637
  • 68799 08/03/01 14:36 R. Michaels Asym Gen Results and Time Asy
  • 68796 08/03/01 14:31 apar Start_of_Run_Parity_1636
  • 68795 08/03/01 14:15 kramerk Rebooted pol3he-mon
  • 68794 08/03/01 14:02 averett new target encoder positions
  • 68793 08/03/01 13:59 A. Deur Raster knobs are flipped
  • 68791 08/03/01 13:41 xfzhu Shift Summary
  • 68788 08/03/01 13:38 A. Deur Raster
  • 68782 08/03/01 13:24 apar Start_of_Run_Parity_1635
  • 68779 08/03/01 13:22 apar Start_of_Run_Parity_1634
  • 68777 08/03/01 13:21 A. Deur DAQ
  • 68774 08/03/01 13:20 apar Start_of_Run_Parity_1633
  • 68771 08/03/01 13:16 averett harp scan/otr
  • 68768 08/03/01 13:14 apar Start_of_Run_Parity_1632
  • 68762 08/03/01 12:42 apar Start_of_Run_Parity_1631
  • 68759 08/03/01 09:53 A. Deur HARP
  • 68757 08/03/01 09:33 apar Start_of_Run_Parity_1630
  • 68753 08/03/01 08:30 R. Michaels re: raster problem ?
  • 68750 08/03/01 08:26 apar Start_of_Run_Parity_1629
  • 68746 08/03/01 08:18 apar Start_of_Run_Parity_1628
  • 68744 08/03/01 08:08 nsavvinv raster problem
  • 68742 08/03/01 08:01 apar Start_of_Run_Parity_1627
  • 68736 08/03/01 07:45 apar Start_of_Run_Parity_1626
  • 68731 08/03/01 07:24 apar Start_of_Run_Parity_1625
  • 68726 08/03/01 07:14 apar Start_of_Run_Parity_1624
  • 68723 08/03/01 06:49 apar Start_of_Run_Parity_1623
  • 68717 08/03/01 06:37 apar Start_of_Run_Parity_1622
  • 68713 08/03/01 06:00 adaq Runsummary_Left_HRS
  • 68712 08/03/01 06:00 adaq Runsummary_Right_HRS
  • 68710 08/03/01 05:58 apar Start_of_Run_Parity_1621
  • 68705 08/03/01 05:53 apar Start_of_Run_Parity_1620
  • 68702 08/03/01 05:51 apar Start_of_Run_Parity_1619
  • 68697 08/03/01 05:40 Reitz harp scan failed
  • 68696 08/03/01 05:36 xfzhu target position problem fixed
  • 68695 08/03/01 05:24 apar Start_of_Run_Parity_1618
  • 68694 08/03/01 05:12 frullani target position problem
  • 68693 08/03/01 04:38 xfzhu Problems with target position.
  • 68692 08/03/01 04:30 xfzhu CW beam
  • 68691 08/03/01 03:18 xfzhu Pulser beam back in hall!
  • 68690 08/03/01 02:43 S.Frullani NMR measurement
  • 68687 08/03/01 01:57 apar Start_of_Run_Parity_1617
  • 68679 08/03/01 01:41 xfzhu Beam is coming
  • 68677 08/03/01 01:02 apar Start_of_Run_Parity_1616
  • 68676 08/03/01 00:58 apar Start_of_Run_Parity_1615
  • 68675 08/03/01 00:54 apar Start_of_Run_Parity_1614
  • 68671 08/03/01 00:28 xfzhu plot for Left arm GC chananel #9: amplitude spectrum
  • Thursday

  • 68668 08/02/01 22:40 apar Start_of_Run_Parity_1613
  • 68666 08/02/01 22:29 Wolfgang Thanks!!
  • 68665 08/02/01 21:54 apar Start_of_Run_Parity_1612
  • 68663 08/02/01 21:08 averett daily summary
  • 68662 08/02/01 20:22 X. Zheng Laser/EPR work during switch-over
  • 68661 08/02/01 18:45 Reitz double pulsing on t1
  • 68660 08/02/01 18:06 Reitz DAQ Synchr reconnected
  • 68658 08/02/01 18:05 apar Start_of_Run_Parity_1611
  • 68652 08/02/01 17:54 apar Start_of_Run_Parity_1610
  • 68647 08/02/01 17:50 apar Start_of_Run_Parity_1609
  • 68642 08/02/01 17:43 apar Start_of_Run_Parity_1608
  • 68625 08/02/01 15:51 xfzhu Left arm GC chananel #9: amplitude spectrum
  • 68624 08/02/01 15:40 xfzhu scaler reading for left arm GC
  • 68623 08/02/01 14:21 R. Michaels Asymmetry Generator tests finished
  • 68620 08/02/01 13:53 Reitz HRSR_S2R_1 TDC peak/ prog width fan out
  • 68619 08/02/01 13:12 xfzhu HV setting for left arm .
  • 68618 08/02/01 13:11 xfzhu scaler reading for left arm GC counter
  • 68616 08/02/01 12:02 apar Start_of_Run_Parity_1607
  • 68612 08/02/01 11:50 gilman,reitz g2 helicity setup
  • 68611 08/02/01 11:44 apar Start_of_Run_Parity_1606
  • 68608 08/02/01 09:45 A. Deur First g2n setting
  • 68607 08/02/01 09:38 R. Michaels helicity electronics problem
  • 68606 08/02/01 09:38 A. Deur quads cycling
  • 68603 08/02/01 09:18 gilman more details on "g2" helicity gating modifications
  • 68602 08/02/01 09:17 A. Deur right Q1
  • 68601 08/02/01 08:14 xfzhu left arm Q2 tripped.
  • 68600 08/02/01 06:00 adaq Runsummary_Left_HRS
  • 68599 08/02/01 06:00 adaq Runsummary_Right_HRS
  • 68596 08/02/01 00:21 apar Start_of_Run_Parity_1605
  • 68591 08/02/01 00:01 apar Start_of_Run_Parity_1604
  • Wednesday

  • 68587 08/01/01 22:03 xfzhu Left arm GC detector with comic ray run2433
  • 68586 08/01/01 21:46 xfzhu Right arm shower counter HV
  • 68585 08/01/01 21:37 xfzhu shower counter with cosmic ray run
  • 68583 08/01/01 21:29 apar Start_of_Run_Parity_1603
  • 68576 08/01/01 21:06 apar Start_of_Run_Parity_1602
  • 68575 08/01/01 20:41 Zolnierczuk tkQA/tkABU scripts
  • 68574 08/01/01 19:38 Zolnierczuk tkQA/tkABU scripts
  • 68573 08/01/01 19:35 Reitz helicity gating change
  • 68567 08/01/01 17:50 A. Deur eP
  • 68554 08/01/01 16:53 apar Start_of_Run_Parity_1601
  • 68553 08/01/01 16:50 apar Start_of_Run_Parity_1600
  • 68552 08/01/01 16:50 R. Michaels mods to parity daq: 10 sec delay
  • 68551 08/01/01 16:44 apar Start_of_Run_Parity_1599
  • 68550 08/01/01 16:41 apar Start_of_Run_Parity_1598
  • 68549 08/01/01 16:31 apar Start_of_Run_Parity_1597
  • 68548 08/01/01 15:53 apar Start_of_Run_Parity_1596
  • 68545 08/01/01 15:45 gilman random test with other helicity modules
  • 68544 08/01/01 15:43 apar Start_of_Run_Parity_1591
  • 68543 08/01/01 15:42 apar Start_of_Run_Parity_1590
  • 68542 08/01/01 15:39 apar Start_of_Run_Parity_1587
  • 68541 08/01/01 15:36 apar Start_of_Run_Parity_1586
  • 68540 08/01/01 15:35 apar Start_of_Run_Parity_1585
  • 68539 08/01/01 15:35 apar Start_of_Run_Parity_1584
  • 68538 08/01/01 15:33 apar Start_of_Run_Parity_1583
  • 68537 08/01/01 15:30 apar Start_of_Run_Parity_1582
  • 68536 08/01/01 15:30 apar Start_of_Run_Parity_1581
  • 68535 08/01/01 15:29 apar Start_of_Run_Parity_1580
  • 68532 08/01/01 15:25 apar Start_of_Run_Parity_1579
  • 68530 08/01/01 15:23 apar Start_of_Run_Parity_1578
  • 68529 08/01/01 15:22 apar Start_of_Run_Parity_1577
  • 68528 08/01/01 15:22 apar Start_of_Run_Parity_1576
  • 68527 08/01/01 15:21 apar Start_of_Run_Parity_1575
  • 68526 08/01/01 14:48 Eugene Ch., Shower #46 and #48
  • 68525 08/01/01 14:28 R. Michaels Asym Test, A1n config, random
  • 68523 08/01/01 14:02 Michaels/Vac fast feedback on Qasy
  • 68521 08/01/01 12:46 apar Start_of_Run_Parity_1574
  • 68518 08/01/01 12:37 R. Michaels Asymmetry Generator Results (II)
  • 68516 08/01/01 12:15 apar Start_of_Run_Parity_1573
  • 68511 08/01/01 12:04 apar Start_of_Run_Parity_1572
  • 68510 08/01/01 12:00 R. Michaels experiment switchover
  • 68509 08/01/01 11:40 Reitz switched to pseudo random
  • 68502 08/01/01 06:00 adaq Runsummary_Left_HRS
  • 68501 08/01/01 06:00 adaq Runsummary_Right_HRS